### v1

# DIVERSE: Deciphering Internet Views on the U.S. Military Through Video Comment Stance Analysis, A Novel Benchmark Dataset for Stance Classification

Abstract
Stance detection of social media text is a key component of downstream tasks involving the identification of groups of users with opposing opinions on contested topics such as vaccination and within arguments. In particular, stance provides an indication of an opinion towards an entity. This paper introduces DIVERSE [^1] The dataset is released at https://doi.org/10.5281/zenodo.10493803, a dataset of over 173,000 YouTube video comments annotated for their stance towards videos of the U.S. military. The stance is annotated through a human-guided, machine-assisted labeling methodology that makes use of weak signals of tone within the sentence as supporting indicators, as opposed to using manual annotations by humans. These weak signals consist of the presence of hate speech and sarcasm, the presence of specific keywords, the sentiment of the text, and the stance inference from two Large Language Models. The weak signals are then consolidated using a data programming model before each comment is annotated with a final stance label. On average, the videos have 200 comments each, and the stance of the comments skews slightly towards the “against” characterization for both the U.S. Army and the videos posted on the channel.
Introduction
At the time of the publication of this article, the United States is facing a military recruitment crisis. For the last few years, the U.S. Military, and the U.S. Army in particular, have struggled to meet recruiting targets to maintain their desired force structure (Brown 2023; Roxana Tiron 2022). Because of the salience of this crisis to the United States’ national well-being (i.e., a professional and ready military is an important component of a healthy state), there has been increased focus from various sectors of government, industry, and academia on discovering why this recruiting crisis is happening and what might be done to address its causes. While there is no agreed-upon reason or set of reasons for why the crisis exists, it is clear that social media plays an important role in military recruiting (Asch 2019; Ng and Cruickshank 2023; Calkins and Asch 2022; Myers 2022; Brown 2023). Thus, it is critical to understand how recruitment messaging is received on social media and how it might be shaping the beliefs of social media users toward something like joining the military. More broadly speaking, the ability to evaluate how social media content is being received by a target audience is crucial for any marketing campaign, whether it be for service in the military or for influencing consumer behavior.
A particular indicator of the reception of social media content is stance. Stance detection, or the determination of an expressed or implied opinion toward a target, remains a primary task for many computational social, social media, and marketing studies. As such, several articles have proposed various methods and datasets for the task of stance detection (Ng and Carley 2022; Allaway and McKeown 2023; Du Bois 2007). However, all existing datasets explore stance as expressed on a single social media platform, Twitter, and only have a single stance per comment (Sobhani, Inkpen, and Zhu 2017; Conforti et al. 2020; Mohammad et al. 2016; Hossain et al. 2020; Kochkina, Liakata, and Zubiaga 2018; Villa-Cox et al. 2020). Furthermore, since stance detection remains a difficult task due to the context-dependent nature of the definition of stance (e.g., the requirement of a target for stance to be defined), as well as varying definitions of stance (Ng and Carley 2022), all of the aforementioned stance datasets rely on human annotation to create the stance labels. This can result in human errors being present in the datasets as well as greatly reduce the size of the datasets to be in the thousands since human annotation requires a lot of resources and time.
To address these shortcomings in prior stance datasets and to aid in understanding the critical issues with the U.S. Army’s recruitment effort, we propose a new stance dataset based on YouTube comments and labeled by human-assisted machine labeling. Our contributions are as follows:
  • A novel benchmark stance dataset of YouTube comments. We present a dataset for the task of stance detection that has several novel differences from previous benchmark datasets. These differences include using comments from a different social media platform (YouTube rather than Twitter), comments having labeled stances toward more than one target (the only other being Election2016 (Sobhani, Inkpen, and Zhu 2017)), the presence of contentious issues, conspiracy theories, and other emotional content, and having more than three times the number of labeled comments over the next largest benchmark stance dataset (the next largest benchmark is wtwt at ${}32,000$ data points (Conforti et al. 2020)).
  • The use of a human-guided, machine-assisted stance-labeling methodology. We adopt recent advances in the creation of large, high-quality labeled datasets (Ratner et al. 2016, 2019), combined with knowledge about the domain and stance, to produce a methodology that can generate quality stance labels at a much larger scale than human-annotation-only methods. This involves making use of weak labels such as the presence of hate speech, the sentiment of the content, the presence of sarcasm, known keywords, and the evaluation of Large Language Models (LLMs) to determine the final stance label.
This dataset is a collection of YouTube comments posted on the videos uploaded by the U.S. Army. This set of comments is annotated for weak labels representing tone (hate speech, sarcasm, and sentiment) before combining with predetermined keywords and the inference of two LLMs to describe the final stance of the comment towards the U.S. Army.
Related Work
Stance Datasets
The task of stance detection has led the community to develop several datasets. Ng and Carley (2022) consolidated several stance detection datasets with social media posts across topics like coronavirus misconceptions, political reactions and tweets revolving around crisis events. Generally, these datasets use three basic labels – favor, against, neither – to express how the sentence views a target (Mohammad et al. 2016). There are datasets that use labels of implicit and explicit support and denial as stances to indicate the overtness of stance expression (Qazvinian et al. 2011). Automated stance detection algorithms have been constructed with these datasets using support vector machines (Elfardy and Diab 2016), logistic regressions (Augenstein, Vlachos, and Bontcheva 2016), neural network models (Fang et al. 2019; Siddiqua, Chy, and Aono 2019) and deep learning models (Küçük and Can 2020). Recently, Large-Language Models have also been used for identifying stance in sentences through direct inference and through (Zhang et al. 2023a; Gatto, Sharif, and Preum 2023).
YouTube Datasets
To date, there have been several research investigations involving YouTube, but few datasets. For example, there are a number of papers investigating influence campaigns (Marcoux et al. 2021; Hussain et al. 2018), conspiracy theories (Liaw et al. 2023), comment toxicity (Obadimu et al. 2019), and even cross-platform use of YouTube videos (Ginossar et al. 2022). Despite the focus on analyzing YouTube data, including comment data, most research projects do not release their datasets (Liaw et al. 2023). Furthermore, to the best of the authors’ knowledge, there are no datasets or investigations of YouTube data for stance detection or classification, particularly concerning the comments section on YouTube videos. It is important to note, however, that there are a number of commercially available social media suites that analyze comments to social media posts on sites like YouTube for things like sentiment and influencer engagement, but none of these tools provide stance detection or classification (Ng and Carley 2022; Rogan et al. 2022).
The Use of Large Language Models for Stance Detection
Recently, researchers have explored the application of Large Language Models (LLMs) in stance detection. Existing studies have primarily focused on ChatGPT, yielding mixed outcomes. For example, Zhang et al. demonstrated improved results on the SemEval2016 benchmark dataset using ChatGPT with instruction-based prompts, outperforming supervised models (Zhang, Ding, and Jing 2022). Other works have also found good, if not necessarily state-of-the-art performance, on the task of stance detection for LLMs (Ziems et al. 2023; Liyanage, Gokani, and Mago 2023; Zhang et al. 2023b). Conversely, Aiyappa et al. raised concerns about potential data contamination from ChatGPT’s extensive training dataset, casting doubt on the reliability of evaluations (Aiyappa et al. 2023). Mets et al. examined ChatGPT’s usability as a zero-shot classifier for stance detection, reporting competitive but ultimately inferior performance compared to supervised models (Mets et al. 2023). Despite these findings, the effectiveness of LLMs in stance classification remains promising, if uncertain, especially without fine-tuning on labeled data or prompt engineering.
Data Programming for Dataset creation
Recently, Data Programming has emerged as a promising approach for expeditiously generating large training datasets. This methodology leverages heuristic rules and weak supervision to programmatically label data at scale (Ratner et al. 2016, 2019). Rather than relying solely on manually annotated datasets, Data Programming enables the automatic creation of training sets by combining various weak sources, such as heuristics, distant supervision, and domain expertise. These weaker, noisy labels can then be combined to produce high-quality labels. This paradigm shift offers an efficient means of tackling data scarcity challenges, especially in scenarios where obtaining labeled data is prohibitively time-consuming or expensive, such as for stance detection. The incorporation of Data Programming into the realm of stance detection holds the potential to create quality labeled datasets with much less human annotation.
Data
In this section, we detail how the dataset was collected, we describe the annotation process, and we comment on the FAIR aspects of the dataset. For this dataset, we labeled each comment for two stances: the stance of the comment toward the video it is commenting on and the stance of the comment toward the U.S. Army. For each stance, we allowed for three possible options: “supports”, “against”, and “neutral”.
Data Collection
This dataset was collected to understand the reception of YouTube users toward the U.S. Army’s marketing material on YouTube. As mentioned in the introduction, social media marketing is a critical aspect of military recruiting, and military recruiting in the United States is undergoing a crisis. Since the U.S. Army maintains an official channel on YouTube, we collected the comments on the videos of that channel to understand how content from the Army is being received and hopefully gain some insight into the causes behind the recruiting crisis. To do this, we first collected comments from all of the videos posted on the U.S. Army’s official YouTube channel (username: @usarmy [^2] https://www.youtube.com/usarmy) using the YouTube Data API [^3] https://developers.google.com/youtube/v3/docs/comments/list. The collection of comments was only from those videos that enabled comments and were accessible via the YouTube API. We collected both comments as well as any replies to comments. Collection of the comments took place from 2 October 2023 to 5 October 2023. In total, we collected 173,013 comments across 858 videos. For each comment, we collected the following information:
  • id: the unique id of a comment
  • comment: the text of the comment
  • author: a unique id for the author of the comment
  • like_count: the likes the comment received
  • published_at: the time and date of the comment’s publication
  • conversation_id: a unique id of the conversation that a comment belongs to. This is also the id of the root comment
  • video_id: the unique id of the video the comment is commenting on
  • name: the title of the video the comment is commenting on
  • description: a text description of the video the comment is commenting on
  • timestamp: the time and date of the publication of the video the comment is commenting on
Data Annotation Methodology
To annotate the data for stance labels, instead of using human annotators, we followed the Data Programming Paradigm (Ratner et al. 2016, 2019). This paradigm produces a final, more accurate label using a series of lesser quality labels, weak labels, which may capture different aspects of the concept being labeled. Such labeling techniques have been employed in labeling happiness with cognitive appraisal dimension (Liu and Jaidka 2023) and negotiation strategies using the presence of reasoning and friendliness (Jaidka, Ahuja, and Ng 2023). We used a combination of weak labels based upon subject matter expertise, domain expertise, machine learning models, and recent advances in the use of LLMs for stance detection. These labels were then combined through a Data Programming model to produce the final labels. Figure 1, depicts the entire data annotation methodology workflow. In the following subsections, we detail key aspects of the workflow, with a special emphasis on the creation of weak labels.
Figure 1: Pipeline of Stance Dataset Creation Methodology
Pre-tuning of Large Language Models
Based on previous research that has shown the possible utility of stance labeling by LLMs, we opted to use LLMs to label comments for their stances. Due to the size of the dataset, we opted to use open-source LLMs which could be run on local hardware. Specifically, we used two different LLMs: Flan-UL2 (Tay et al. 2022), which is an encoder-decoder architecture and Mistral-7B (Jiang et al. 2023), which is a decoder-only architecture. Based on past research (Ziems et al. 2023), we used the Flan-UL2 model in a zero-shot setting (i.e., we did not perform any fine-tuning on the model). For the Mistral-7B model, we fine-tuned the model on a collection of benchmark stance datasets, including Semeval2016 (Mohammad et al. 2016), Election2016 (Sobhani, Inkpen, and Zhu 2017), srq (Villa-Cox et al. 2020), wtwt (Conforti et al. 2020), phemerumors (Kochkina, Liakata, and Zubiaga 2018), and covid-lies (Hossain et al. 2020). Based on the research in (Ng and Carley 2022), which found that the inclusion of more benchmark datasets into training improved the generalizability of the trained model, hence we trained on all of the datasets. For the training of the Mistral LLM, we used a LoRA adapter (Hu et al. 2021) with parameters of $\alpha=1$, $r=16$, $dropout=0.1$, and no bias term. The model was trained for two epochs using an AdamW optimizer (Loshchilov and Hutter 2017), a learning rate of $41e-$ 4, and a linear learning rate scheduler with 50 warm-up steps. To feed the data to the model during training, we adopted a prompt, with two variations in the task portion of the prompt, that included context about stance classification, as has been done in zero-shot prompting for stance in other works. The following was the training prompt:
“‘Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a certain, specified target. Analyze the following social media statement and determine its stance towards the provided [dataset entity].
Respond with a single word: “for”, “against”,“neutral”.
: {entity}
statement: {statement to classify}
stance: ”’
Where the [dataset entity] was varied depending on the benchmark dataset (e.g., for Election2016, it was ‘politician’, while for phemerumors it was ‘rumor’). For the other variation of the prompt, the “Analyze…” sentence was replaced with a question-based tasking in the form of: “What is the stance of the following social media statement towards the provided [dataset entity]”. For the srq dataset, we changed the prompt slightly to incorporate the post that the reply is replying to as the task in that dataset is to label the stance of the reply to the post, versus the stance of the post about a topic or person. The loss was determined by whether the model could output the tokens for the correct stance, and only those tokens. The training of the model was done on a local computer running Ubuntu Linux with a 96-core processor and three A6000 GPUs. Overall, this resulted in three, different LLMs for labeling the comments: one off-the-shelf UL2 model and two, fine-tuned Mistral models.
Weak Labels
To label the comments for stance, we first needed to produce noisy, weak labels. To do this we leveraged dataset-specific traits, domain expertise and LLMs. For example, in this dataset, the presence of hate speech or sarcasm frequently indicates the comment being against the video and by extension, the user that uploads it, which is the US Army. Therefore, the presence of these weak signals can be combined to indicate the final label. Finally, it is important to note for any given weak label, it need not produce a label for every comment in the dataset. Some of the weak labels simply do not apply to some of the comments, in which case they have an “abstain” response, which is distinct from the possible stance labels, and are not used in the computation of the final stance labels.
Hate Speech
Hate Speech is defined as speech to incite violence or hate, and typically has specific targets like ethnic groups or religion (Fortuna and Nunes 2018). In the case of this dataset and stance, we noticed that many of the comments against both the posted videos and the U.S. Army more generally featured aspects of hate speech, like antisemitic and misogynistic content. Several machine learning algorithms ranging from logistic regression to neural network models have been constructed to perform automatic hate speech detection (Fortuna and Nunes 2018). For this weak label, we used the dehatebert model from HuggingFace [^4] Hate-speech-CNERG/dehatebert-mono-english (Aluru et al. 2020), a multi-lingual fine-tuned BERT model. This model was used as is without any fine-tuning. For each input text, this model returns a binary label of “hate” or “no_hate”, and we considered the ”hate” response as an “against” stance label.
Sarcasm
Sarcasm is a literary device to convey one’s intention in an implicit manner (Chaudhari and Chandankhede 2017). As with hate speech, we found that commenters on these YouTube videos often use sarcasm as a way of attacking the video or the entity behind the video. We used a multilingual sarcasm detector555helinivan/multilingual-sarcasm-detector which trains a BERT-based sarcasm detector from newspapers text in English, Dutch and Italian. For each input text, this model returns a label of “sarcasm” or “no_sarcasm”, and we considered the “sarcasm” response as an “against” stance label
Sentiment
Sentiment analysis provides an understanding of the positivity or negativity of public sentiment towards a topic (Ibrohim, Bosco, and Basile 2023). In the context of stance detection and classification, while stance is distinct from sentiment, the two often overlap; a negative sentiment often indicates a stance against a target and vice-versa for a positive sentiment. We used a multilingual-cased sentiment analysis model666lxyuan/distilbert-base-multilingual-cased-sentiments-student that performs sentiment-analysis using a teacher-student architecture with the DistillBERT model. For each input text, this model returns a score between 0 and 1 for three possible sentiment labels of “positive”, “negative”, and “neutral”. We considered a “positive” sentiment greater than 0.75 as a “supports” stance and a “negative” sentiment greater than 0.75 as a “against” stance and any other value as a “neutral” stance.
Use of Keywords
We manually determined a set of keywords and phrases that upon present, would represent either one of the stance labels. For each comment, we identified if any words from the set of keywords were present and assigned the corresponding stance to this weak label. For example, keywords and phrases like “hooah”, “awesome”, and “god bless” were associated with a “supports” stance, while “woke”, “murder”, and “propaganda” were associated with an “against” stance.
Stance from Large Language Models
Based on the recent work demonstrating some effectiveness of LLMs in the stance classification task, we also directly labeled the comments for their stances using LLMs. We used two different prompts, one that only included the comment, and one that included the comment that a comment was replying to, if it was replying to a comment, with two different variations on each prompt. The labeling prompt without incorporating the reply structure is:
“‘Analyze the following YouTube comment to a video posted by the U.S. Army named “{title}” and determine its stance towards the provided entity. Respond with a single word: “for”, “against”, “neutral”, or “unrelated”. Only return the stance as a single word, and no other text.
entity: {entity}
comment: {comment to classify}
stance:
while the labeling prompt that did incorporate the reply structure is:
“‘Analyze the following YouTube comment to a video posted by the U.S. Army named “{title}” and determine its stance towards the provided entity. If the statement is a reply to another YouTube comment, that YouTube comment is listed below. Respond with a single word: “for”, “against”, “neutral”, or “unrelated”. Only return the stance as a single word, and no other text.
entity: {entity}
previous comment: {previous_comment}
comment: {comment to classify}
stance:
As with the training prompt, the variations for each prompt came from changing the task part of the prompt (i.e., the “Analyze…”) sentence to a question-based task prompt. These variations were only used with the Mistral model fine-tuned by the question prompt Thus, each comment was labeled for a stance by an LLM six different times: once with the reply prompt and once without the reply prompt, for the UL2 model, the stance-tuned Mistral model, and the stance-tuned Mistral model with the question-task prompt. Since we are labeling the comment for its stance toward the U.S. Army as well as the video it is commenting on, this results in a total of 12 LLM labeling runs across the entire dataset.
Use of Other Stances
Finally, in the case of labeling the stance of the comment toward its respective video, we also used the stance of the comment toward the U.S. Army as a weak label for this stance. Generally speaking, if a comment was against the U.S. Army, it would also be against a video posted by the U.S. Army, and the use of complementary stances has been shown to improve performance in the stance detection task (Sobhani, Inkpen, and Zhu 2017; Zhang et al. 2023b). Thus, we used the final stance toward the U.S. Army produced by Data Programming on the aforementioned weak labels as a weak label for the stance toward the video. It is important to note that this relationship does not work the other way; we found several examples where commentators would express a supportive stance toward the U.S. Army but would be very against a particular video.
Determining Stance
Having created the weak labels for the comments, we then use the Data Programming methodology in the form of the snorkel Python package [^7] https://snorkel.readthedocs.io/en/master/index.html, which uses a unified generative model that takes into account partial coverage of the weak labels as well as correlations between the weak labels to produce the final labels. We refer the reader to (Ratner et al. 2016) and (Ratner et al. 2019) for the specifics on the function, how it is trained, and how it produces the final labels. For the final labels, we allow the stance to take the form of “supports”, “against”, and “neutral” for both the stance toward the video and the stance toward the U.S. Army.
FAIR Aspects of Dataset
Findability & Accessibility
We provide public access to our dataset and source code at the indexed-dataset sharing service Zenodo, https://doi.org/10.5281/zenodo.10493803 to facilitate replication and use of this dataset. A meta-data file is also provided to briefly describe each of the fields. In particular, we also provide the weak labels as part of the dataset as well.
Interoperability & Reusability
The dataset is provided in a generic .csv format which supports reusability across several data analysis applications. We also provide the code used to generate the dataset for further reusability. This code includes the code used to produce the weak labels as well as the code to create the final labels from the weak labels.
Ethical Considerations
In the collection of this dataset, we only collected data provided by the YouTube API and did not collect any private profiles or videos. We anonymized the author IDs, both in the author field of the data as well as in the comments section if a comment were a reply to another comment, and we did not perform any author-level analysis.
Dataset Properties
In this section, we outline some properties of the data as well as the stance labels to understand the dataset.
Data Properties
The dataset consists of 858 videos published from 1 June 2010 through 5 October 2023 (final date of collection), all from the U.S. Army’s official YouTube channel. Figure 2 depicts the distribution of the videos on the channel over time
Figure 2: Distribution of videos published on the U.S. Army’s Official Channel over time
Generally speaking, the channel has published an increasing number of videos over time, with some occasional periods of fewer videos. Table 1 summarizes the statistics for the comments to each video. Most videos within the data set only have a single comment, with the number of comments per video being highly skewed and some videos having a large number of comments (the most commented-on video, ‘Soldier wakes up and chooses “HOOAH!”’, has 9.4% of all of the comments by itself). When it comes to the timing of the comments themselves, we find the overwhelming have occurred recently in the channel’s timeline (since 2022). Figure 3, depicts the distribution of comments, across all videos, by month and year.
Statistics of comments per video Value
Mean 201.65
Standard Deviation 873.19
Minimum 1
25th Percentile 5
50th Percentile (Median) 27
75th Percentile 77
Maximum 16289
Mode 1
Table 1: Per Video Comment Statistics
Figure 3: Distribution of videos published on the U.S. Army’s Official Channel over time
From the plot of when comments are published, there are significant spikes in comments in August, September, and October of 2022 and again in May and June of 2023. In the case of the first spike, the most commented-on video was ‘Soldier wakes up and chooses “HOOAH!”’, while the second spike in the Spring of 2023 has ‘Words of wisdom from a #WWII veteran’ the most commented-on video. For the first video, it seems the commentary is divisively split about what the Army does or should do in the world, while in the second case the comments seem to both positively reminisce about the Army in World War II and lament about its current state. Turning to the comments themselves, we find that most comments are relatively short, like many social media posts, but a few comments are much longer at over 1,000 words for the comment. Table 2 has the word count statistics for the comments.
Word-level Statistics of Comments Value
Mean 16.77
Standard Deviation 32.67
Minimum 1.00
25th Percentile 4.00
50th Percentile (Median) 9.00
75th Percentile 18.00
Maximum 1851.00
Table 2: Word-level Statistics of the Comments
Label Properties
Turning now to stance labels for the dataset, we find that the labels skew slightly toward the “against” characterization for both the U.S. Army and the videos being posted on the channel. The following table, 3 summarizes the label counts for each of the stances.
Stance U.S. Army Video
against 63676 (34.87%) 67405 (36.92%)
neutral 49153 (26.91%) 49310 (27.01%)
supports 60183 (32.99%) 56297 (30.88%)
Table 3: Labels Counts (and percentages of the total number) for Each of the Stance Targets
Interestingly, from the label counts we also observe both a relatively even distribution of labels across the stances but also the “neutral” stance being the minority label. In most other benchmark stance datasets, there is a strong label skew, particularly for datasets like covid-lies (Hossain et al. 2020) and phemerumors (Kochkina, Liakata, and Zubiaga 2018), the “neutral” stance tends to be the majority stance label. Since this is a first-of-its-kind dataset for both the targets as well as the data domain (i.e., YouTube) it’s not clear if the targets tend to inspire stronger stance-taking behavior in comments or if the domain of YouTube causes stronger stance-taking behavior. Regardless, upon manual inspection of some of the labels, we do find that these stance labels are accurate and that there is a lot of strong stance-taking within the comments.
Turning to the distribution of stance labels over time, we find that each of the three classes of stance generally trend together over time. Figure 4 displays the distribution of the different classes of stance toward the U.S. Army for all of the comments (the stances toward the videos follow a very similar trend). From this figure, we can observe that the two prominent periods of comments do differ somewhat in their compositions of stance labels. In the first peak of comments, centering in September of 2022, we see more “against” stance labels than “supports” or “neutral”. Whereas in the second spike of comments around May-June of 2023, we see more “supports” comments than “against” or “neutral”. Thus, overall, there is a moderate balance among the stances expressed in the comments toward the videos, but this balance does shift over time and between videos.
Figure 4: Distribution of videos published on the U.S. Army’s Official Channel over time
To better understand these differences in stances, we also looked at the videos with the most positive, neutral and negative receptions. For videos with more than 100 comments, We found the video, “Honoring Our Fallen Heroes - Taps” had the highest percentage of supportive comments at 76.5% of the comments supporting the video. On the opposite end was the video “What does the Army of the future look like?” which features a short interview with former Sergeant Major Grinston on the future of the U.S. Army, which had 66.9% of comments that were against the video. And, the most neutrally received video was “Army #PopQuiz - What training is this?” which received 66.5% of its comments as being neutral toward the video, many of them answering the question posed in the video.
Finally, since this dataset was labeled with more than one stance, we can also look at the relationships between the two stances. Figure 5, displays a confusion matrix between the stance toward the U.S. Army and the stance toward the videos. Not unsurprisingly, we find these stances closely resemble each other. It is, however, worth pointing out that we do see a significant number of comments that are supportive of the video, but neutral toward the Army as well as against or neutral to the video, but supportive of the Army. In the first case, from some manual inspection, these are often comments that give a positive word about the video and nothing else. For example, a comment like “that video was awesome!” fits into this category. For, the second category these comments are often comments that express some frustration or anger at a particular video, but are otherwise supportive of the Army. For example, a comment like “What is this bs? The Army is better than this…” Overall, the two stance labels closely resemble each other in this dataset, but there are some notable exceptions where they differ, which highlights some of the complexity of stance within this domain and more generally.
Figure 5: Confusion Matrix of Stance Toward the U.S. Army and Stance Toward the Video
Leave-One-Out Stance Detection
To further validate the stance labels produced by the proposed Data Programming method, we performed a leave-one-out, out-of-domain experiment referenced from (Ng and Carley 2022). We trained a model on the stance labels and then evaluated that model on the other benchmark datasets. In our test, we trained a cased BERT model [^8] https://huggingface.co/bert-base-cased (Devlin et al. 2018) on only the stance labels from the DISMISS dataset and then evaluated the model on all of the benchmark datasets. The model was trained for 3 epochs using a learning rate of $41e-$ 4 with linear decay, a batch size of 32, and was trained to classify text into three different labels. The text that was input to the model was given as:
“target: {target of stance} [SEP] comment: {comment}”
In this test, if the proposed data labels are capturing the stances correctly, then we should see similar performance on the benchmark datasets when predicting stance, as was seen in (Ng and Carley 2022). Table 4 has the out-of-domain test results. The results are evaluated using the Macro F1 score as a suitable evaluation method due to the almost equal distribution of stance labels.
On average, the model trained on stances only from this dataset achieves a macro F1 score of 0.3043. This is similar to the average accuracy of 0.303 in a past work that performed leave-one-out stance detection using multiple BERT-based classifiers (Ng and Carley 2022). This therefore suggests that the stance labels proposed in this paper are valid stance labels.
Dataset Macro F1 Score
semeval2016 0.3517
election2016 0.3439
covid-lies 0.3020
phemerumors 0.2295
wtwt 0.3018
srq 0.2971
Average 0.3043
Table 4: Macro F1 Scores for Model Applied to Different Datasets
Discussion
Your text is well-written, but I’ve made some minor corrections for clarity and consistency. Here’s the revised version:
Public opinion towards a country’s army is crucial, as it signals the level of trust in the army and the willingness of citizens to join. A positive opinion facilitates army recruitment. Social media engagement plays a pivotal role in shaping this opinion. Therefore, this paper introduces a dataset that supports such understanding and associated analyses, comprising comments on the U.S. Army’s official YouTube videos annotated for their stance towards the Army and the video content itself.
To construct this dataset, we employed the Data Programming paradigm, which leverages weak labels for automated labeling. In this paradigm, humans transition from direct annotation to machine-assisted annotation, allowing scalability to larger datasets. For this specific dataset, we utilized subject matter expertise, domain knowledge, and recent advancements in using LLMs for stance determination to create weak labels. Specifically, we used weak labels such as the presence of hate speech, sarcasm, and sentiment as indicators for certain stances. This transformation converted the challenging, ambiguously defined problem of determining and labeling stance into a more well-defined problem for specific aspects or stances. These could then be aggregated through Data Programming to produce comprehensive stance labeling. For instance, the presence of hate speech indicates a comment against the video and, by extension, against the user who uploaded it (e.g., the Army).
We further enhanced these weak labels with contributions from LLMs. Currently, the effectiveness of LLMs in tasks like stance labeling remains uncertain. While recent research shows promise for using LLMs, especially when combined with prompt engineering, for stance detection and labeling, the generalizability of these methods and their compatibility with all LLMs is unclear. Through the Data Programming label, we treat these LLM-derived labels as noisy yet valuable, enhancing the creation of a stance-labeled dataset. The combination of dataset-specific weak signals and the linguistic understanding power of LLMs produces high-quality labels for stance in a more scalable manner compared to purely manual or machine-generated labels.
Finally, our Leave-One-Out experiment produces a standard macro F1 score of around 0.34, a score that is similar to past work. This suggests that the stance labels produced by the proposed methodology are indeed valid. This result also highlights findings from past work that trained stance detection models do not necessarily generalize to other data domains (e.g., YouTube versus Twitter, or targets of videos versus political figures, etc.). Thus, we suggest training a stance classification model using multiple benchmark datasets so the model can better recognize the nuances of the stances and targets involved. The low F1 score also indicates the difficulty of the stance classification task, whereby language models do struggle to differentiate the nuances in opinions towards a target entity, despite having helper annotations like hate and sentiment to provide indications of stances. This, therefore, opens up avenues for exploration of improving the stance classification methods, especially on out-of-domain tasks, to provide a more accurate label towards an entity and therefore more accurate downstream insights.
Potential Applications
The dataset proposed in this work has several potential uses. First, this dataset presents a versatile resource for research focusing on the analysis of stance in YouTube video comments. Beyond its primary application in training and evaluating new stance detection techniques and models, the dataset introduces distinctive traits not present in current benchmarks. Notably, it originates from YouTube, offering a departure from the prevalent Twitter-based datasets. Furthermore, the dataset features multiple stance targets and exhibits a substantial degree of stance-taking behavior and controversial content (i.e., a high proportion of “supports” and “against” stances). The dataset provides sequential information about comments, allowing investigations into the dynamics of stance with respect to the stance expressed in a previous comment.
Given the dataset’s proximity to a politically and culturally divisive topic — the U.S. Military — it encompasses a spectrum of online behaviors, including the propagation of conspiracy theories, trolling, hate speech, and misinformation. As such, it serves as a valuable resource for studies on misinformation and disinformation.
Moreover, the dataset’s temporal collection allows researchers to explore dynamic aspects of stance-taking behavior over time. This temporal dimension provides insights into the evolution of discussions and sentiments within the context of the U.S. Army’s YouTube channel. In summary, the dataset offers a rich foundation for various studies, ranging from advancing stance detection methodologies to delving into the nuanced dynamics of online interactions related to a prominent and controversial subject.
Limitations
Our dataset is limited to the comments focusing on the US Army, and only those we could collect. Not all of the videos posted to the channel allowed comments. We hope to expand our dataset to include the comments towards videos posted by armies around the world, to evaluate public opinion towards the army. In addition, our dataset is limited to the stance annotation of local, relatively small LLMs due to the size of the dataset and computational constraints. In the future, we hope to expand the dataset annotation to different types of LLMs with varying prompting schemes to better understand the stance annotation capabilities of LLMs.
Conclusion
Understanding the stance expressed in YouTube comments towards an entity is essential for gaining insights into public opinion regarding the authors of the videos. In this context, we introduce the DIVERSE dataset, a dataset that captures the stance towards the US Army. This dataset is formulated from comments gathered from the U.S. Army’s YouTube videos. We constructed the final stance variable utilizing weak labels, including the identification of hate speech and sarcasm, and incorporated stance inferences from Large Language Models. By employing a data programming model, we amalgamated these diverse labels to generate the final stance annotations. We anticipate that this dataset will prove invaluable to researchers investigating public sentiments and stances toward military entities.
Acknowledgments
The research for this paper was supported in part by the Center for Informed Democracy and Social-cybersecurity (IDeaS) and the Center for Computational Analysis of Social and Organizational Systems (CASOS) at Carnegie Mellon University. This work was also conducted within the Cognitive Security Research Lab at the Army Cyber Institute. The views and conclusions are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of Defense, the U.S. Army, or the U.S. Government.
Appendix A Appendix 1: Ethics Paper Checklist
  1. 1.
    For most authors…
    1. (a)
      Would answering this research question advance science without violating social contracts, such as violating privacy norms, perpetuating unfair profiling, exacerbating the socio-economic divide, or implying disrespect to societies or cultures? Yes
    2. (b)
      Do your main claims in the abstract and introduction accurately reflect the paper’s contributions and scope? Yes
    3. (c)
      Do you clarify how the proposed methodological approach is appropriate for the claims made? Yes
    4. (d)
      Do you clarify what are possible artifacts in the data used, given population-specific distributions? We discuss the limitations of the dataset as well as give basic statistics on the dataset to clarify population-level distributions.
    5. (e)
      Did you describe the limitations of your work? Limitations are described in the limitations subsection in the discussion section.
    6. (f)
      Did you discuss any potential negative societal impacts of your work? Yes
    7. (g)
      Did you discuss any potential misuse of your work? Yes
    8. (h)
      Did you describe steps taken to prevent or mitigate potential negative outcomes of the research, such as data and model documentation, data anonymization, responsible release, access control, and the reproducibility of findings? Yes. We follow standard anonymization procedures for the social media data
    9. (i)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? Yes
  2. 2.
    Additionally, if your study involves hypotheses testing…
    1. (a)
      Did you clearly state the assumptions underlying all theoretical results? NA
    2. (b)
      Have you provided justifications for all theoretical results? NA
    3. (c)
      Did you discuss competing hypotheses or theories that might challenge or complement your theoretical results? NA
    4. (d)
      Have you considered alternative mechanisms or explanations that might account for the same outcomes observed in your study? NA
    5. (e)
      Did you address potential biases or limitations in your theoretical framework? NA
    6. (f)
      Have you related your theoretical results to the existing literature in social science? NA
    7. (g)
      Did you discuss the implications of your theoretical results for policy, practice, or further research in the social science domain? NA
  3. 3.
    Additionally, if you are including theoretical proofs…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? NA
    2. (b)
      Did you include complete proofs of all theoretical results? NA
  4. 4.
    Additionally, if you ran machine learning experiments…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? We include all of the code in a URL
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? Yes.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? NA
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? Yes.
    5. (e)
      Do you justify how the proposed evaluation is sufficient and appropriate to the claims made? Yes.
    6. (f)
      Do you discuss what is “the cost“ of misclassification and fault (in)tolerance? Yes.
  5. 5.
    Additionally, if you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? NA
    2. (b)
      Did you mention the license of the assets? NA
    3. (c)
      Did you include any new assets in the supplemental material or as a URL? NA
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? NA. All data is publicly available data. We were also asked to analyze this data by the organization that owns the YouTube channel from which the data comes from.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? Yes. We mention the data contains various types of conspiracy theories, hate speech ad misinformation. We also anonymize the data to remove user names to and within comments
    6. (f)
      If you are curating or releasing new datasets, did you discuss how you intend to make your datasets FAIR (see fair)? Yes.
    7. (g)
      If you are curating or releasing new datasets, did you create a Datasheet for the Dataset (see gebru2021datasheets)? Yes.
  6. 6.
    Additionally, if you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots? NA
    2. (b)
      Did you describe any potential participant risks, with mentions of Institutional Review Board (IRB) approvals? NA
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? NA
    4. (d)
      Did you discuss how data is stored, shared, and deidentified? Yes

### v2

@@
-DIVERSE: Deciphering Internet Views on the U.S. Military Through Video Comment Stance Analysis, A Novel Benchmark Dataset for Stance Classification
+DIVERSE: A Dataset of YouTube Video Comment Stances with a Data Programming Model
 Abstract
-Stance detection of social media text is a key component of downstream tasks involving the identification of groups of users with opposing opinions on contested topics such as vaccination and within arguments.
-In particular, stance provides an indication of an opinion towards an entity.
-This paper introduces DIVERSE [^1] The dataset is released at https://doi.org/10.5281/zenodo.10493803, a dataset of over 173,000 YouTube video comments annotated for their stance towards videos of the U.S.
-military.
-The stance is annotated through a human-guided, machine-assisted labeling methodology that makes use of weak signals of tone within the sentence as supporting indicators, as opposed to using manual annotations by humans.
-These weak signals consist of the presence of hate speech and sarcasm, the presence of specific keywords, the sentiment of the text, and the stance inference from two Large Language Models.
-The weak signals are then consolidated using a data programming model before each comment is annotated with a final stance label.
-On average, the videos have 200 comments each, and the stance of the comments skews slightly towards the “against” characterization for both the U.S.
-Army and the videos posted on the channel.
-Introduction
-At the time of the publication of this article, the United States is facing a military recruitment crisis.
-For the last few years, the U.S.
-Military, and the U.S.
-Army in particular, have struggled to meet recruiting targets to maintain their desired force structure (Brown 2023; Roxana Tiron 2022).
-Because of the salience of this crisis to the United States’ national well-being (i.e., a professional and ready military is an important component of a healthy state), there has been increased focus from various sectors of government, industry, and academia on discovering why this recruiting crisis is happening and what might be done to address its causes.
-While there is no agreed-upon reason or set of reasons for why the crisis exists, it is clear that social media plays an important role in military recruiting (Asch 2019; Ng and Cruickshank 2023; Calkins and Asch 2022; Myers 2022; Brown 2023).
-Thus, it is critical to understand how recruitment messaging is received on social media and how it might be shaping the beliefs of social media users toward something like joining the military.
-More broadly speaking, the ability to evaluate how social media content is being received by a target audience is crucial for any marketing campaign, whether it be for service in the military or for influencing consumer behavior.
+Stance detection of social media text is a key component of many real-world applications like evaluating marketing campaigns, evaluating political policies or candidates, or evaluating information environments.
+However, creating automatic stance labeling systems requires the manual annotation of stances, which is both tedious and resource-intensive.
+This paper introduces an stance labeling method that makes use of weak signals of sentence tone, then consolidating these signals with a Data Programmingmodel for the final stance label.
+In a time of international conflict, understanding the public opinion towards the country’s military is crucial for recruitment.
+We present DIVERSE, a dataset involve stances towards YouTube videos of the US military [^1] Datasetavailableathttps://doi.org/10.5281/zenodo.10493803.
+On average, the videos have 200 comments each, and the stances skew slightly towards the “against” characterization for both the US army and the video.
+Index Terms:
+stance detection, social media, large language models, ensemble learning
+I Introduction
+Today’s world is rife with armed conflict: the Russia-Ukraine war, the Israel-Hamas conflict, the Yemeni Civil War, Somali Civil War… and so forth.
+The US is involved in a number of these conflicts, and also has the mammoth task of defending her own.
+During conflict, it is crucial to know the stance of the people towards its own defending forces, as a gauge of the support the forces have from the public.
+However, US military is facing a recruitment crisis.
+For the last few years, the US Army have struggled to meet recruiting targets to maintain their desired force structure (Brown, 2023).
+Social media plays an important role in discerning stances (AlDayel and Magdy, 2021), and in recruitment (Ng and Cruickshank, 2023).
+Thus, it is critical to understand how messages from the military is received on social media, for it may be used to shape beliefs of users.
+More broadly speaking, the ability to evaluate how social media content is being received by a target audience is crucial for any marketing campaign, be it a war recruiting effort or a organizational recruitment effort or for influencing consumer behavior.
 A particular indicator of the reception of social media content is stance.
-Stance detection, or the determination of an expressed or implied opinion toward a target, remains a primary task for many computational social, social media, and marketing studies.
-As such, several articles have proposed various methods and datasets for the task of stance detection (Ng and Carley 2022; Allaway and McKeown 2023; Du Bois 2007).
-However, all existing datasets explore stance as expressed on a single social media platform, Twitter, and only have a single stance per comment (Sobhani, Inkpen, and Zhu 2017; Conforti et al.
-2020; Mohammad et al.
-2016; Hossain et al.
-2020; Kochkina, Liakata, and Zubiaga 2018; Villa-Cox et al.
-2020).
-Furthermore, since stance detection remains a difficult task due to the context-dependent nature of the definition of stance (e.g., the requirement of a target for stance to be defined), as well as varying definitions of stance (Ng and Carley 2022), all of the aforementioned stance datasets rely on human annotation to create the stance labels.
+Stance detection, or the determination of an expressed or implied opinion toward a target, remains a primary task for many computational social studies.
+Several articles have proposed methods and datasets for the task of stance detection (Allaway and McKeown, 2023; Du Bois, 2007).
+However, existing datasets explore stance as expressed on a single social media platform, mainly X, and only have a single stance per comment (Sobhani et al., 2017; Conforti et al., 2020; Mohammad et al., 2016; Hossain et al., 2020; Kochkina et al., 2018; Villa-Cox et al., 2020).
+Since stance detection remains a difficult task due to the context-dependent nature of the definition of stance (e.g., the requirement of a target for stance to be defined), along with varying definitions of stance (Ng and Carley, 2022), all of the aforementioned stance datasets rely on human annotation to create the stance labels.
 This can result in human errors being present in the datasets as well as greatly reduce the size of the datasets to be in the thousands since human annotation requires a lot of resources and time.
-To address these shortcomings in prior stance datasets and to aid in understanding the critical issues with the U.S.
-Army’s recruitment effort, we propose a new stance dataset based on YouTube comments and labeled by human-assisted machine labeling.
-Our contributions are as follows:
+To address these shortcomings in prior stance dataset, we propose a new stance dataset based on YouTube comments and labeled with machine labeling through the data programming paradigm of weak supervision.
+This dataset is a collection of YouTube comments posted on videos uploaded by the US Army [^2] Offical US Army YouTube Channel: https://www.youtube.com/@usarmy.
+The comments are annotated for weak labels representing tone (hate speech, sarcasm, sentiment), inference of two LLMs (stance towards video, stance towards army), then combining these labels to produce the final stance labels of the comment towards the US Army.
+With this dataset, we develop a automated and systematic approach to labeling stances, aiding in understanding critical issues with the public opinion towards the army and its recruitment efforts.
+Our contributions of this paper are as follows:
+  • The use of an data-programming based stance-labeling methodology.
+We combine recent advances for creating large, high-quality labeled datasets Ratner et al.
+(2016, 2019) with contextual knowledge to generate quality stance labels.
+This involves making use of weak labels such as the presence of hate speech, the sentiment of the content, the presence of sarcasm, known keywords, and the evaluation of Large Language Models (LLMs), which are combined using weak supervision to determine the final stance label.
   • A novel benchmark stance dataset of YouTube comments.
 We present a dataset for the task of stance detection that has several novel differences from previous benchmark datasets.
-These differences include using comments from a different social media platform (YouTube rather than Twitter), comments having labeled stances toward more than one target (the only other being Election2016 (Sobhani, Inkpen, and Zhu 2017)), the presence of contentious issues, conspiracy theories, and other emotional content, and having more than three times the number of labeled comments over the next largest benchmark stance dataset (the next largest benchmark is wtwt at ${}32,000$ data points (Conforti et al.
-2020)).
-  • The use of a human-guided, machine-assisted stance-labeling methodology.
-We adopt recent advances in the creation of large, high-quality labeled datasets (Ratner et al.
-2016, 2019), combined with knowledge about the domain and stance, to produce a methodology that can generate quality stance labels at a much larger scale than human-annotation-only methods.
-This involves making use of weak labels such as the presence of hate speech, the sentiment of the content, the presence of sarcasm, known keywords, and the evaluation of Large Language Models (LLMs) to determine the final stance label.
-This dataset is a collection of YouTube comments posted on the videos uploaded by the U.S.
-Army.
-This set of comments is annotated for weak labels representing tone (hate speech, sarcasm, and sentiment) before combining with predetermined keywords and the inference of two LLMs to describe the final stance of the comment towards the U.S.
-Army.
-Related Work
+While most stance datasets use posts from Twitter, ours provides posts from the YouTube platform instead, contributing to the breadth of datasets.
+Another difference is that our dataset labels stances towards more than one target in a comment (the only other being Sobhani et al.
+(2017)), and provides labels to weak signals including: the presence of contentious issues, conspiracy theories, and other emotional content.
+Finally, this dataset has more than three times the number of labeled comments over the current largest benchmark dataset (Conforti et al.
+(2020) at ${}32,000$ data points vs 173,013 data points).
+II Related Work
 Stance Datasets
-The task of stance detection has led the community to develop several datasets.
-Ng and Carley (2022) consolidated several stance detection datasets with social media posts across topics like coronavirus misconceptions, political reactions and tweets revolving around crisis events.
-Generally, these datasets use three basic labels – favor, against, neither – to express how the sentence views a target (Mohammad et al.
-2016).
-There are datasets that use labels of implicit and explicit support and denial as stances to indicate the overtness of stance expression (Qazvinian et al.
-2011).
-Automated stance detection algorithms have been constructed with these datasets using support vector machines (Elfardy and Diab 2016), logistic regressions (Augenstein, Vlachos, and Bontcheva 2016), neural network models (Fang et al.
-2019; Siddiqua, Chy, and Aono 2019) and deep learning models (Küçük and Can 2020).
-Recently, Large-Language Models have also been used for identifying stance in sentences through direct inference and through (Zhang et al.
-2023a; Gatto, Sharif, and Preum 2023).
+The task of stance detection has led to the development of several datasets.
+Ng and Carley (2022) consolidated several stance detection datasets with social media posts across multiple topics.
+These datasets use three basic labels – favor, against, neither – to express how the sentence views a target (Mohammad et al., 2016).
+There are datasets that use labels of implicit and explicit support and denial to indicate the overtness of stance expression (Qazvinian et al., 2011).
+Automated stance detection algorithms have been constructed with these datasets using support vector machines (Elfardy and Diab, 2016), logistic regressions (Augenstein et al., 2016), neural network models (Fang et al., 2019; Siddiqua et al., 2019) and deep learning models (Küçük and Can, 2020).
+Recently, Large-Language Models have also been used for identifying stance in sentences (Zhang et al., 2023a; Gatto et al., 2023).
 YouTube Datasets
 To date, there have been several research investigations involving YouTube, but few datasets.
-For example, there are a number of papers investigating influence campaigns (Marcoux et al.
-2021; Hussain et al.
-2018), conspiracy theories (Liaw et al.
-2023), comment toxicity (Obadimu et al.
-2019), and even cross-platform use of YouTube videos (Ginossar et al.
-2022).
-Despite the focus on analyzing YouTube data, including comment data, most research projects do not release their datasets (Liaw et al.
-2023).
-Furthermore, to the best of the authors’ knowledge, there are no datasets or investigations of YouTube data for stance detection or classification, particularly concerning the comments section on YouTube videos.
-It is important to note, however, that there are a number of commercially available social media suites that analyze comments to social media posts on sites like YouTube for things like sentiment and influencer engagement, but none of these tools provide stance detection or classification (Ng and Carley 2022; Rogan et al.
-2022).
-The Use of Large Language Models for Stance Detection
+For example, there are a number of papers investigating influence campaigns (Marcoux et al., 2021; Hussain et al., 2018), conspiracy theories (Liaw et al., 2023), comment toxicity (Obadimu et al., 2019), and even cross-platform use of YouTube videos (Ginossar et al., 2022).
+Despite the focus on analyzing YouTube data, including comment data, most research projects do not release their datasets (Liaw et al., 2023).
+Furthermore, to the best of the authors’ knowledge, there are no datasets or investigations of YouTube data for stance detection, particularly concerning the comments section on YouTube videos.
+It is important to note, however, that there are a number of commercially available social media suites to analyze comments to posts on platforms for indicators like sentiment and influencer engagement, but none of these tools provide stance detection (Ng and Carley, 2022; Rogan et al., 2022).
+Large Language Models for Stance Detection
 Recently, researchers have explored the application of Large Language Models (LLMs) in stance detection.
-Existing studies have primarily focused on ChatGPT, yielding mixed outcomes.
-For example, Zhang et al.
-demonstrated improved results on the SemEval2016 benchmark dataset using ChatGPT with instruction-based prompts, outperforming supervised models (Zhang, Ding, and Jing 2022).
-Other works have also found good, if not necessarily state-of-the-art performance, on the task of stance detection for LLMs (Ziems et al.
-2023; Liyanage, Gokani, and Mago 2023; Zhang et al.
-2023b).
+Existing studies primarily focused on ChatGPT, yielding mixed outcomes.
+Zhang et al.
+(2022) demonstrated improved results on the SemEval2016 benchmark dataset using ChatGPT with instruction-based prompts.
+Other works have also found good, if not necessarily state-of-the-art performance, on the task of stance detection for LLMs (Ziems et al., 2023; Liyanage et al., 2023; Zhang et al., 2023b).
 Conversely, Aiyappa et al.
-raised concerns about potential data contamination from ChatGPT’s extensive training dataset, casting doubt on the reliability of evaluations (Aiyappa et al.
-2023).
+(2023) raised concerns about potential data contamination from ChatGPT’s extensive training dataset, casting doubt on the reliability of evaluations.
 Mets et al.
-examined ChatGPT’s usability as a zero-shot classifier for stance detection, reporting competitive but ultimately inferior performance compared to supervised models (Mets et al.
-2023).
+(2023) examined ChatGPT’s usability as a zero-shot classifier for stance detection, reporting competitive but ultimately inferior performance compared to supervised models.
 Despite these findings, the effectiveness of LLMs in stance classification remains promising, if uncertain, especially without fine-tuning on labeled data or prompt engineering.
-Data Programming for Dataset creation
-Recently, Data Programming has emerged as a promising approach for expeditiously generating large training datasets.
-This methodology leverages heuristic rules and weak supervision to programmatically label data at scale (Ratner et al.
-2016, 2019).
-Rather than relying solely on manually annotated datasets, Data Programming enables the automatic creation of training sets by combining various weak sources, such as heuristics, distant supervision, and domain expertise.
+Weak Supervision for Dataset creation
+Weak Supervision has emerged as a promising approach for expeditiously generating large training datasets.
+This methodology leverages heuristic rules and weak supervision to programmatically label data at scale (Ratner et al., 2016, 2019).
+Rather than relying solely on manually annotated datasets, weak supervision enables the automatic creation of training sets by combining various weak sources, derived from heuristics, distant supervision, and domain expertise.
 These weaker, noisy labels can then be combined to produce high-quality labels.
-This paradigm shift offers an efficient means of tackling data scarcity challenges, especially in scenarios where obtaining labeled data is prohibitively time-consuming or expensive, such as for stance detection.
-The incorporation of Data Programming into the realm of stance detection holds the potential to create quality labeled datasets with much less human annotation.
-Data
-In this section, we detail how the dataset was collected, we describe the annotation process, and we comment on the FAIR aspects of the dataset.
-For this dataset, we labeled each comment for two stances: the stance of the comment toward the video it is commenting on and the stance of the comment toward the U.S.
-Army.
-For each stance, we allowed for three possible options: “supports”, “against”, and “neutral”.
-Data Collection
-This dataset was collected to understand the reception of YouTube users toward the U.S.
-Army’s marketing material on YouTube.
-As mentioned in the introduction, social media marketing is a critical aspect of military recruiting, and military recruiting in the United States is undergoing a crisis.
-Since the U.S.
-Army maintains an official channel on YouTube, we collected the comments on the videos of that channel to understand how content from the Army is being received and hopefully gain some insight into the causes behind the recruiting crisis.
-To do this, we first collected comments from all of the videos posted on the U.S.
-Army’s official YouTube channel (username: @usarmy [^2] https://www.youtube.com/usarmy) using the YouTube Data API [^3] https://developers.google.com/youtube/v3/docs/comments/list.
-The collection of comments was only from those videos that enabled comments and were accessible via the YouTube API.
-We collected both comments as well as any replies to comments.
-Collection of the comments took place from 2 October 2023 to 5 October 2023.
-In total, we collected 173,013 comments across 858 videos.
+This paradigm shift offers an efficient means of tackling data scarcity challenges, especially in scenarios where obtaining labeled data is time-consuming or expensive.
+The incorporation of weak supervision into the realm of stance detection holds the potential to create quality labeled datasets with much less human annotation.
+III Dataset Creation
+This dataset contains comments of YouTube comments, each labeled for a set of two final stances: (a) the stance of the comment towards its source video, and (b) the stance of the comment towards the US Army.
+There are three possible options for stances: “supports”, “against”, and “neutral”.
+This dataset also contains the weak labels used to fold into the final stance through weak supervision.
+III-A Data Collection
+This dataset was collected to understand the reception of YouTube users toward the US Army’s marketing material on YouTube, thereby understanding their support/opposition towards the military.
+We collected comments on the videos of the US Army’s official YouTube channel (username: @usarmy [^3] https://www.youtube.com/usarmy) using the YouTube Data API [^4] https://developers.google.com/youtube/v3/docs/comments/list.
+This returned both comments as well as replies to comments.
+Collection took place from 2 to 5 October 2023, where we retrieved a total of 173,013 comments across 858 videos.
 For each comment, we collected the following information:
@@
   • timestamp: the time and date of the publication of the video the comment is commenting on
-Data Annotation Methodology
-To annotate the data for stance labels, instead of using human annotators, we followed the Data Programming Paradigm (Ratner et al.
-2016, 2019).
+III-B Data Annotation Methodology
+Instead of using human annotators for stance labeling, we used weak supervision Ratner et al.
+(2016, 2019).
 This paradigm produces a final, more accurate label using a series of lesser quality labels, weak labels, which may capture different aspects of the concept being labeled.
-Such labeling techniques have been employed in labeling happiness with cognitive appraisal dimension (Liu and Jaidka 2023) and negotiation strategies using the presence of reasoning and friendliness (Jaidka, Ahuja, and Ng 2023).
+Such labeling techniques have been employed in labeling happiness with cognitive appraisal dimension Liu and Jaidka (2023) and negotiation strategies using the presence of reasoning and friendliness Jaidka et al.
+(2023).
 We used a combination of weak labels based upon subject matter expertise, domain expertise, machine learning models, and recent advances in the use of LLMs for stance detection.
-These labels were then combined through a Data Programming model to produce the final labels.
+These labels were then combined through a weak supervision model to produce the final labels.
 Figure 1, depicts the entire data annotation methodology workflow.
-In the following subsections, we detail key aspects of the workflow, with a special emphasis on the creation of weak labels.
-Figure 1: Pipeline of Stance Dataset Creation Methodology
-Pre-tuning of Large Language Models
-Based on previous research that has shown the possible utility of stance labeling by LLMs, we opted to use LLMs to label comments for their stances.
+We detail key aspects of the workflow in the following sections.
+Figure 1: Pipeline of stance labeling methodology note: change LLM box to 2 boxes not 3
+III-C Annotating Weak Labels
+To label the comments for stance, we first produced weak labels.
+To do this we leveraged dataset-specific traits, domain expertise and LLMs.
+For example, in this dataset, the presence of hate speech or sarcasm frequently indicates the comment being against the video and by extension, the user that uploads it, which is the US Army.
+Therefore, the presence of these weak signals can be combined to indicate the final label.
+It is important to note for any given weak label, it need not produce a label for every comment in the dataset.
+Some of the weak labels simply do not apply to some of the comments, in which case they have an “abstain” response, which is distinct from the possible stance labels, and are not used in the computation of the final stance labels.
+III-C1 Hate Speech
+Hate Speech is defined as speech to incite violence or hate, and typically has specific targets like ethnic groups or religion Fortuna and Nunes (2018).
+In the case of this dataset and stance, we noticed that many of the comments against both the posted videos and the US Army more generally featured aspects of hate speech, like antisemitic and misogynistic content.
+Several machine learning algorithms ranging from logistic regression to neural network models have been constructed to perform automatic hate speech detection Fortuna and Nunes (2018).
+For this weak label, we used the dehatebert model from HuggingFace [^5] Hate-speech-CNERG/dehatebert-mono-english Aluru et al.
+(2020), a multi-lingual fine-tuned BERT model.
+This model was used as is without any fine-tuning.
+For each input text, this model returns a binary label of “hate” or “no_hate”, and we considered the “hate” response as an “against” stance label.
+III-C2 Sarcasm
+Sarcasm is a literary device to convey one’s intention in an implicit manner Chaudhari and Chandankhede (2017).
+As with hate speech, we found that commenters on these YouTube videos often use sarcasm as a way of attacking the video or the entity behind the video.
+We used a multilingual sarcasm detector666helinivan/multilingual-sarcasm-detector which trains a BERT-based sarcasm detector from newspapers text in English, Dutch and Italian.
+For each input text, this model returns a label of “sarcasm” or “no_sarcasm”, and we considered the “sarcasm” response as an “against” stance label
+III-C3 Sentiment
+Sentiment analysis provides an understanding of the positivity or negativity of public sentiment towards a topic Ibrohim et al.
+(2023).
+In the context of stance detection and classification, while stance is distinct from sentiment, the two often overlap; a negative sentiment often indicates a stance against a target and vice-versa for a positive sentiment.
+We used a multilingual-cased sentiment analysis model777lxyuan/distilbert-base-multilingual-cased-sentiments-student that performs sentiment-analysis using a teacher-student architecture with the DistillBERT model.
+For each input text, this model returns a score between 0 and 1 for three possible sentiment labels of “positive”, “negative”, and “neutral”.
+We considered a “positive” sentiment greater than 0.75 as a “supports” stance and a “negative” sentiment greater than 0.75 as a “against” stance and any other value as a “neutral” stance.
+III-C4 Use of Keywords
+We manually determined a set of keywords and phrases that upon present, would represent either one of the stance labels.
+For each comment, we identified if any words from the set of keywords were present and assigned the corresponding stance to this weak label.
+For example, keywords and phrases like “hooah”, “awesome”, and “god bless” were associated with a “supports” stance, while “woke”, “murder”, and “propaganda” were associated with an “against” stance.
+III-C5 Stance from Large Language Models
+Based on the recent work demonstrating effectiveness of LLMs in the stance classification task (Ng et al., 2024), we also directly labeled stances using LLMs.
+Pre-tuning of LLMs
 Due to the size of the dataset, we opted to use open-source LLMs which could be run on local hardware.
-Specifically, we used two different LLMs: Flan-UL2 (Tay et al.
-2022), which is an encoder-decoder architecture and Mistral-7B (Jiang et al.
-2023), which is a decoder-only architecture.
-Based on past research (Ziems et al.
-2023), we used the Flan-UL2 model in a zero-shot setting (i.e., we did not perform any fine-tuning on the model).
-For the Mistral-7B model, we fine-tuned the model on a collection of benchmark stance datasets, including Semeval2016 (Mohammad et al.
-2016), Election2016 (Sobhani, Inkpen, and Zhu 2017), srq (Villa-Cox et al.
-2020), wtwt (Conforti et al.
-2020), phemerumors (Kochkina, Liakata, and Zubiaga 2018), and covid-lies (Hossain et al.
-2020).
-Based on the research in (Ng and Carley 2022), which found that the inclusion of more benchmark datasets into training improved the generalizability of the trained model, hence we trained on all of the datasets.
-For the training of the Mistral LLM, we used a LoRA adapter (Hu et al.
-2021) with parameters of $\alpha=1$, $r=16$, $dropout=0.1$, and no bias term.
-The model was trained for two epochs using an AdamW optimizer (Loshchilov and Hutter 2017), a learning rate of $41e-$ 4, and a linear learning rate scheduler with 50 warm-up steps.
+Specifically, we used two different LLMs: (a) Flan-UL2 (Tay et al., 2022), which is an encoder-decoder architecture and (b) Mistral-7B (Jiang et al., 2023), which is a decoder-only architecture.
+We used the Flan-UL2 model in a zero-shot setting (i.e., we did not perform any fine-tuning on the model) (Ziems et al., 2023).
+For the Mistral-7B model, we fine-tuned the model on a collection of benchmark stance datasets, including Semeval2016 (Mohammad et al., 2016), Election2016 (Sobhani et al., 2017), srq (Villa-Cox et al., 2020), wtwt (Conforti et al., 2020), phemerumors (Kochkina et al., 2018), and covid-lies (Hossain et al., 2020).
+Based on the research in (Ng and Carley, 2022), which found that the inclusion of more benchmark datasets into training improved the generalizability of the trained model, hence we trained on all of the datasets.
+For the training of the Mistral LLM, we used a LoRA adapter (Hu et al., 2021) with parameters of $\alpha=1$, $r=16$, $dropout=0.1$, and no bias term.
+The model was trained for two epochs using an AdamW optimizer (Loshchilov and Hutter, 2017), a learning rate of $41e-$ 4, and a linear learning rate scheduler with 50 warm-up steps.
 To feed the data to the model during training, we adopted a prompt, with two variations in the task portion of the prompt, that included context about stance classification, as has been done in zero-shot prompting for stance in other works.
@@
 Respond with a single word: “for”, “against”,“neutral”.
-: {entity}
+[dataset entity]: {entity}
 statement: {statement to classify}
@@
 Overall, this resulted in three, different LLMs for labeling the comments: one off-the-shelf UL2 model and two, fine-tuned Mistral models.
-Weak Labels
-To label the comments for stance, we first needed to produce noisy, weak labels.
-To do this we leveraged dataset-specific traits, domain expertise and LLMs.
-For example, in this dataset, the presence of hate speech or sarcasm frequently indicates the comment being against the video and by extension, the user that uploads it, which is the US Army.
-Therefore, the presence of these weak signals can be combined to indicate the final label.
-Finally, it is important to note for any given weak label, it need not produce a label for every comment in the dataset.
-Some of the weak labels simply do not apply to some of the comments, in which case they have an “abstain” response, which is distinct from the possible stance labels, and are not used in the computation of the final stance labels.
-Hate Speech
-Hate Speech is defined as speech to incite violence or hate, and typically has specific targets like ethnic groups or religion (Fortuna and Nunes 2018).
-In the case of this dataset and stance, we noticed that many of the comments against both the posted videos and the U.S.
-Army more generally featured aspects of hate speech, like antisemitic and misogynistic content.
-Several machine learning algorithms ranging from logistic regression to neural network models have been constructed to perform automatic hate speech detection (Fortuna and Nunes 2018).
-For this weak label, we used the dehatebert model from HuggingFace [^4] Hate-speech-CNERG/dehatebert-mono-english (Aluru et al.
-2020), a multi-lingual fine-tuned BERT model.
-This model was used as is without any fine-tuning.
-For each input text, this model returns a binary label of “hate” or “no_hate”, and we considered the ”hate” response as an “against” stance label.
-Sarcasm
-Sarcasm is a literary device to convey one’s intention in an implicit manner (Chaudhari and Chandankhede 2017).
-As with hate speech, we found that commenters on these YouTube videos often use sarcasm as a way of attacking the video or the entity behind the video.
-We used a multilingual sarcasm detector555helinivan/multilingual-sarcasm-detector which trains a BERT-based sarcasm detector from newspapers text in English, Dutch and Italian.
-For each input text, this model returns a label of “sarcasm” or “no_sarcasm”, and we considered the “sarcasm” response as an “against” stance label
-Sentiment
-Sentiment analysis provides an understanding of the positivity or negativity of public sentiment towards a topic (Ibrohim, Bosco, and Basile 2023).
-In the context of stance detection and classification, while stance is distinct from sentiment, the two often overlap; a negative sentiment often indicates a stance against a target and vice-versa for a positive sentiment.
-We used a multilingual-cased sentiment analysis model666lxyuan/distilbert-base-multilingual-cased-sentiments-student that performs sentiment-analysis using a teacher-student architecture with the DistillBERT model.
-For each input text, this model returns a score between 0 and 1 for three possible sentiment labels of “positive”, “negative”, and “neutral”.
-We considered a “positive” sentiment greater than 0.75 as a “supports” stance and a “negative” sentiment greater than 0.75 as a “against” stance and any other value as a “neutral” stance.
-Use of Keywords
-We manually determined a set of keywords and phrases that upon present, would represent either one of the stance labels.
-For each comment, we identified if any words from the set of keywords were present and assigned the corresponding stance to this weak label.
-For example, keywords and phrases like “hooah”, “awesome”, and “god bless” were associated with a “supports” stance, while “woke”, “murder”, and “propaganda” were associated with an “against” stance.
-Stance from Large Language Models
-Based on the recent work demonstrating some effectiveness of LLMs in the stance classification task, we also directly labeled the comments for their stances using LLMs.
-We used two different prompts, one that only included the comment, and one that included the comment that a comment was replying to, if it was replying to a comment, with two different variations on each prompt.
+Analyzing stances with LLMs
+In analyzing the comments for stances using LLMs, we used two different prompts, one that only included the comment, and one that included the comment that a comment was replying to, if it was replying to a comment, with two different variations on each prompt.
 The labeling prompt without incorporating the reply structure is:
@@
 These variations were only used with the Mistral model fine-tuned by the question prompt Thus, each comment was labeled for a stance by an LLM six different times: once with the reply prompt and once without the reply prompt, for the UL2 model, the stance-tuned Mistral model, and the stance-tuned Mistral model with the question-task prompt.
-Since we are labeling the comment for its stance toward the U.S.
-Army as well as the video it is commenting on, this results in a total of 12 LLM labeling runs across the entire dataset.
-Use of Other Stances
-Finally, in the case of labeling the stance of the comment toward its respective video, we also used the stance of the comment toward the U.S.
-Army as a weak label for this stance.
-Generally speaking, if a comment was against the U.S.
-Army, it would also be against a video posted by the U.S.
-Army, and the use of complementary stances has been shown to improve performance in the stance detection task (Sobhani, Inkpen, and Zhu 2017; Zhang et al.
-2023b).
-Thus, we used the final stance toward the U.S.
-Army produced by Data Programming on the aforementioned weak labels as a weak label for the stance toward the video.
-It is important to note that this relationship does not work the other way; we found several examples where commentators would express a supportive stance toward the U.S.
-Army but would be very against a particular video.
-Determining Stance
-Having created the weak labels for the comments, we then use the Data Programming methodology in the form of the snorkel Python package [^7] https://snorkel.readthedocs.io/en/master/index.html, which uses a unified generative model that takes into account partial coverage of the weak labels as well as correlations between the weak labels to produce the final labels.
-We refer the reader to (Ratner et al.
-2016) and (Ratner et al.
-2019) for the specifics on the function, how it is trained, and how it produces the final labels.
-For the final labels, we allow the stance to take the form of “supports”, “against”, and “neutral” for both the stance toward the video and the stance toward the U.S.
-Army.
-FAIR Aspects of Dataset
-Findability & Accessibility
-We provide public access to our dataset and source code at the indexed-dataset sharing service Zenodo, https://doi.org/10.5281/zenodo.10493803 to facilitate replication and use of this dataset.
-A meta-data file is also provided to briefly describe each of the fields.
-In particular, we also provide the weak labels as part of the dataset as well.
-Interoperability & Reusability
-The dataset is provided in a generic .csv format which supports reusability across several data analysis applications.
-We also provide the code used to generate the dataset for further reusability.
-This code includes the code used to produce the weak labels as well as the code to create the final labels from the weak labels.
-Ethical Considerations
-In the collection of this dataset, we only collected data provided by the YouTube API and did not collect any private profiles or videos.
-We anonymized the author IDs, both in the author field of the data as well as in the comments section if a comment were a reply to another comment, and we did not perform any author-level analysis.
-Dataset Properties
-In this section, we outline some properties of the data as well as the stance labels to understand the dataset.
-Data Properties
-The dataset consists of 858 videos published from 1 June 2010 through 5 October 2023 (final date of collection), all from the U.S.
-Army’s official YouTube channel.
-Figure 2 depicts the distribution of the videos on the channel over time
-Figure 2: Distribution of videos published on the U.S. Army’s Official Channel over time
+Since we are labeling the comment for its stance toward the US Army as well as the video it is commenting on, this results in a total of 12 LLM labeling runs across the entire dataset.
+III-C6 Use of Analogous Stances
+Finally, in the case of labeling the stance of the comment toward its respective video, we also used an analogous stance for a weak label.
+We used the stance of the comment toward the US Army.
+Generally speaking, if a comment was against the US Army, it would also be against a video posted by the US Army, and the use of complementary stances has been shown to improve performance in the stance detection task Sobhani et al.
+(2017); Zhang et al.
+(2023b).
+Thus, we used the final stance toward the US Army produced by weak supervision on the aforementioned weak labels as a weak label for the stance toward the video.
+It is important to note that this relationship does not work the other way; we found several examples where commentators would express a supportive stance toward the US Army but would be very against a particular video.
+III-D Final Stance Determination
+Having created the weak labels for the comments, we then use the weak supervision methodology in the form of the snorkel Python package [^8] https://snorkel.readthedocs.io/en/master/index.html, which uses a unified generative model that takes into account partial coverage of the weak labels as well as correlations between the weak labels to produce the final labels.
+We refer the reader to Ratner et al.
+(2016) and Ratner et al.
+(2019) for the specifics on the function, how it is trained, and how it produces the final labels.
+For the final labels, we allow the stance to take the form of “supports”, “against”, and “neutral” for both the stance toward the video and the stance toward the US Army.
+III-E Human evaluation
+After machine annotation, we had four individuals annotate how many?
+what percentage of total?
+comments.
+For each comment, the authors independently annotated the stance towards the US Army and the stance towards the video.
+The manual annotations were then analyzed for inter-annotator agreement, and the machine annotations were compared against the manual ones.
+IV Data Properties and Analysis
+This section describes some exploratory data analysis performed on this dataset.
+The dataset consists of 858 videos published from 1 June 2010 through 5 October 2023 (final date of collection) on the US Army’s official YouTube Channel.
+IV-A Dataset Properties
+Distribution of videos over time
 Generally speaking, the channel has published an increasing number of videos over time, with some occasional periods of fewer videos.
-Table 1 summarizes the statistics for the comments to each video.
+Figure 3 depicts the distribution of all the videos on the channel over time.
+Table II summarizes the statistics for the comments to each video.
 Most videos within the data set only have a single comment, with the number of comments per video being highly skewed and some videos having a large number of comments (the most commented-on video, ‘Soldier wakes up and chooses “HOOAH!”’, has 9.4% of all of the comments by itself).
 When it comes to the timing of the comments themselves, we find the overwhelming have occurred recently in the channel’s timeline (since 2022).
-Figure 3, depicts the distribution of comments, across all videos, by month and year.
+Figure 2: Distribution of comments over time
+From Figure 2, the plot of comments over time, there are significant spikes in comments in August, September, and October of 2022 and again in May and June of 2023.
+In the case of the first spike, the most commented-on video was ‘Soldier wakes up and chooses “HOOAH!”’, while the second spike in the Spring of 2023 has ‘Words of wisdom from a #WWII veteran’ the most commented-on video.
+For the first video, it seems the commentary is divisively split about what the Army does or should do in the world, while in the second case the comments seem to both positively reminisce about the Army in World War II and lament about its current state.
+Turning to the comments themselves, we find that most comments are relatively short, like many social media posts, but a few comments are much longer at over 1,000 words for the comment.
+Table I has the word count statistics for the comments.
+Word-level Statistics of Comments Value
+Mean 16.77
+Standard Deviation 32.67
+Minimum 1.00
+25th Percentile 4.00
+50th Percentile (Median) 9.00
+75th Percentile 18.00
+Maximum 1851.00
+TABLE I: Word-level Statistics of the Comments
 Statistics of comments per video Value
@@
 Mode 1
-Table 1: Per Video Comment Statistics
-Figure 3: Distribution of videos published on the U.S. Army’s Official Channel over time
-From the plot of when comments are published, there are significant spikes in comments in August, September, and October of 2022 and again in May and June of 2023.
-In the case of the first spike, the most commented-on video was ‘Soldier wakes up and chooses “HOOAH!”’, while the second spike in the Spring of 2023 has ‘Words of wisdom from a #WWII veteran’ the most commented-on video.
-For the first video, it seems the commentary is divisively split about what the Army does or should do in the world, while in the second case the comments seem to both positively reminisce about the Army in World War II and lament about its current state.
-Turning to the comments themselves, we find that most comments are relatively short, like many social media posts, but a few comments are much longer at over 1,000 words for the comment.
-Table 2 has the word count statistics for the comments.
-Word-level Statistics of Comments Value
-Mean 16.77
-Standard Deviation 32.67
-Minimum 1.00
-25th Percentile 4.00
-50th Percentile (Median) 9.00
-75th Percentile 18.00
-Maximum 1851.00
-Table 2: Word-level Statistics of the Comments
-Label Properties
-Turning now to stance labels for the dataset, we find that the labels skew slightly toward the “against” characterization for both the U.S.
-Army and the videos being posted on the channel.
-The following table, 3 summarizes the label counts for each of the stances.
+TABLE II: Per Video Comment Statistics
+Figure 3: Distribution of videos published over time
+IV-B Stance Label Properties
+Turning now to stance labels produced through our methodology for the dataset, we find that the labels skew slightly toward the “against” characterization for both the US Army and the videos being posted on the channel.
+Table III summarizes the label counts for each of the stances.
+Interestingly, from the label counts we also observe both a relatively even distribution of labels across the stances but also the “neutral” stance being the minority label.
+In most other benchmark stance datasets, there is a strong label skew, particularly for datasets like Hossain et al.
+(2020) and Kochkina et al.
+(2018), the “neutral” stance tends to be the majority stance label.
+Since this is a first-of-its-kind dataset for both the targets as well as the data domain (i.e., YouTube) it’s not clear if the targets tend to inspire stronger stance-taking behavior in comments or if the domain of YouTube causes stronger stance-taking behavior.
+Regardless, upon manual inspection of some of the labels, we do find that these stance labels are accurate and that there is a lot of strong stance-taking within the comments.
 Stance U.S. Army Video
@@
 supports 60183 (32.99%) 56297 (30.88%)
-Table 3: Labels Counts (and percentages of the total number) for Each of the Stance Targets
-Interestingly, from the label counts we also observe both a relatively even distribution of labels across the stances but also the “neutral” stance being the minority label.
-In most other benchmark stance datasets, there is a strong label skew, particularly for datasets like covid-lies (Hossain et al.
-2020) and phemerumors (Kochkina, Liakata, and Zubiaga 2018), the “neutral” stance tends to be the majority stance label.
-Since this is a first-of-its-kind dataset for both the targets as well as the data domain (i.e., YouTube) it’s not clear if the targets tend to inspire stronger stance-taking behavior in comments or if the domain of YouTube causes stronger stance-taking behavior.
-Regardless, upon manual inspection of some of the labels, we do find that these stance labels are accurate and that there is a lot of strong stance-taking within the comments.
+TABLE III: Labels Counts (and percentages of the total number) for Each of the Stance Targets
 Turning to the distribution of stance labels over time, we find that each of the three classes of stance generally trend together over time.
-Figure 4 displays the distribution of the different classes of stance toward the U.S.
-Army for all of the comments (the stances toward the videos follow a very similar trend).
-From this figure, we can observe that the two prominent periods of comments do differ somewhat in their compositions of stance labels.
+Figure 4 displays the distribution of the different classes of stance toward the US Army for all of the comments.
+The different stances toward the videos follow a very similar trend.
+We further observe that the two prominent periods of comments do differ somewhat in their compositions of stance labels.
 In the first peak of comments, centering in September of 2022, we see more “against” stance labels than “supports” or “neutral”.
@@
 Figure 4: Distribution of videos published on the U.S. Army’s Official Channel over time
-To better understand these differences in stances, we also looked at the videos with the most positive, neutral and negative receptions.
-For videos with more than 100 comments, We found the video, “Honoring Our Fallen Heroes - Taps” had the highest percentage of supportive comments at 76.5% of the comments supporting the video.
-On the opposite end was the video “What does the Army of the future look like?” which features a short interview with former Sergeant Major Grinston on the future of the U.S.
-Army, which had 66.9% of comments that were against the video.
+To better understand these differences in stances, we looked at the videos with the most positive, neutral and negative receptions.
+Among the videos with more than 100 comments, we found the video, “Honoring Our Fallen Heroes - Taps” had the highest percentage of supportive comments at 76.5% of the comments supporting the video.
+On the opposite end was the video “What does the Army of the future look like?” which features a short interview with former Sergeant Major Grinston on the future of the US Army.
+The video had 66.9% of comments that were against it.
 And, the most neutrally received video was “Army #PopQuiz - What training is this?” which received 66.5% of its comments as being neutral toward the video, many of them answering the question posed in the video.
@@
 Figure 5: Confusion Matrix of Stance Toward the U.S. Army and Stance Toward the Video
-Leave-One-Out Stance Detection
-To further validate the stance labels produced by the proposed Data Programming method, we performed a leave-one-out, out-of-domain experiment referenced from (Ng and Carley 2022).
+IV-C Comparison with human evaluation
+The inter-annotator agreement was Iain, results here. Comparing the machine-based annotations with the human-based annotations, we get an accuracy of how many. elaborate on accuracy
+IV-D Leave-One-Out Stance Detection
+To further validate the stance labels produced by the proposed weak supervision method, we performed a leave-one-out, out-of-domain experiment referenced from (Ng and Carley, 2022).
 We trained a model on the stance labels and then evaluated that model on the other benchmark datasets.
-In our test, we trained a cased BERT model [^8] https://huggingface.co/bert-base-cased (Devlin et al.
-2018) on only the stance labels from the DISMISS dataset and then evaluated the model on all of the benchmark datasets.
+In our test, we trained a cased BERT model [^9] https://huggingface.co/bert-base-cased (Devlin et al., 2018) on only the stance labels from the DISMISS dataset and then evaluated the model on all of the benchmark datasets.
 The model was trained for 3 epochs using a learning rate of $41e-$ 4 with linear decay, a batch size of 32, and was trained to classify text into three different labels.
@@
 “target: {target of stance} [SEP] comment: {comment}”
-In this test, if the proposed data labels are capturing the stances correctly, then we should see similar performance on the benchmark datasets when predicting stance, as was seen in (Ng and Carley 2022).
-Table 4 has the out-of-domain test results.
+In this test, if the proposed data labels are capturing the stances correctly, then we should see similar performance on the benchmark datasets when predicting stance, as was seen in Ng and Carley (2022).
+Table IV has the out-of-domain test results.
 The results are evaluated using the Macro F1 score as a suitable evaluation method due to the almost equal distribution of stance labels.
 On average, the model trained on stances only from this dataset achieves a macro F1 score of 0.3043.
-This is similar to the average accuracy of 0.303 in a past work that performed leave-one-out stance detection using multiple BERT-based classifiers (Ng and Carley 2022).
+This is similar to the average accuracy of 0.303 in a past work that performed leave-one-out stance detection using multiple BERT-based classifiers Ng and Carley (2022).
 This therefore suggests that the stance labels proposed in this paper are valid stance labels.
@@
 Average 0.3043
-Table 4: Macro F1 Scores for Model Applied to Different Datasets
-Discussion
-Your text is well-written, but I’ve made some minor corrections for clarity and consistency. Here’s the revised version:
-Public opinion towards a country’s army is crucial, as it signals the level of trust in the army and the willingness of citizens to join.
-A positive opinion facilitates army recruitment.
+TABLE IV: Macro F1 Scores for Model Applied to Different Datasets
+V Discussion
+Public opinion towards a country’s army is crucial, as it signals the level of trust in the army and the willingness of citizens to join the forces.
+A positive opinion facilitates army recruitment and support, especially in times of crisis.
 Social media engagement plays a pivotal role in shaping this opinion.
-Therefore, this paper introduces a dataset that supports such understanding and associated analyses, comprising comments on the U.S.
-Army’s official YouTube videos annotated for their stance towards the Army and the video content itself.
-To construct this dataset, we employed the Data Programming paradigm, which leverages weak labels for automated labeling.
-In this paradigm, humans transition from direct annotation to machine-assisted annotation, allowing scalability to larger datasets.
-For this specific dataset, we utilized subject matter expertise, domain knowledge, and recent advancements in using LLMs for stance determination to create weak labels.
+This paper introduces a dataset that supports such understanding and associated analyses, comprising comments on the US Army’s official YouTube videos annotated for their stance towards the Army and the video content itself.
+To construct this dataset, we employed weak supervision, which leverages weak labels for automated labeling.
+This paradigm provides a transition from direct human annotation to machine-assisted annotation, allowing scalability to larger datasets.
+For our specific dataset, we utilized subject matter expertise, domain knowledge, and recent advancements in using LLMs for stance determination to create weak labels.
 Specifically, we used weak labels such as the presence of hate speech, sarcasm, and sentiment as indicators for certain stances.
 This transformation converted the challenging, ambiguously defined problem of determining and labeling stance into a more well-defined problem for specific aspects or stances.
-These could then be aggregated through Data Programming to produce comprehensive stance labeling.
+These could then be aggregated through weak supervision to produce comprehensive stance labeling.
 For instance, the presence of hate speech indicates a comment against the video and, by extension, against the user who uploaded it (e.g., the Army).
@@
 While recent research shows promise for using LLMs, especially when combined with prompt engineering, for stance detection and labeling, the generalizability of these methods and their compatibility with all LLMs is unclear.
-Through the Data Programming label, we treat these LLM-derived labels as noisy yet valuable, enhancing the creation of a stance-labeled dataset.
+Through weak supervision, we treat these LLM-derived labels as noisy yet valuable, enhancing the creation of a stance-labeled dataset.
 The combination of dataset-specific weak signals and the linguistic understanding power of LLMs produces high-quality labels for stance in a more scalable manner compared to purely manual or machine-generated labels.
@@
 The dataset proposed in this work has several potential uses.
-First, this dataset presents a versatile resource for research focusing on the analysis of stance in YouTube video comments.
+First, this dataset presents a versatile resource for research on the analysis of stance in YouTube video comments.
 Beyond its primary application in training and evaluating new stance detection techniques and models, the dataset introduces distinctive traits not present in current benchmarks.
-Notably, it originates from YouTube, offering a departure from the prevalent Twitter-based datasets.
+Notably, it originates from YouTube, offering a departure from the prevalent X-based datasets.
 Furthermore, the dataset features multiple stance targets and exhibits a substantial degree of stance-taking behavior and controversial content (i.e., a high proportion of “supports” and “against” stances).
 The dataset provides sequential information about comments, allowing investigations into the dynamics of stance with respect to the stance expressed in a previous comment.
-Given the dataset’s proximity to a politically and culturally divisive topic — the U.S.
-Military — it encompasses a spectrum of online behaviors, including the propagation of conspiracy theories, trolling, hate speech, and misinformation.
+Given the dataset’s proximity to a politically and culturally divisive topic — the military — it encompasses a spectrum of online behaviors, including the propagation of conspiracy theories, trolling, hate speech, and misinformation.
 As such, it serves as a valuable resource for studies on misinformation and disinformation.
-Moreover, the dataset’s temporal collection allows researchers to explore dynamic aspects of stance-taking behavior over time.
-This temporal dimension provides insights into the evolution of discussions and sentiments within the context of the U.S.
-Army’s YouTube channel.
+Finally, the dataset’s temporal collection allows researchers to explore dynamic aspects of stance-taking behavior over time.
+This temporal dimension provides insights into the evolution of discussions and sentiments within the context of the US Army’s YouTube channel.
 In summary, the dataset offers a rich foundation for various studies, ranging from advancing stance detection methodologies to delving into the nuanced dynamics of online interactions related to a prominent and controversial subject.
-Limitations
-Our dataset is limited to the comments focusing on the US Army, and only those we could collect.
-Not all of the videos posted to the channel allowed comments.
-We hope to expand our dataset to include the comments towards videos posted by armies around the world, to evaluate public opinion towards the army.
-In addition, our dataset is limited to the stance annotation of local, relatively small LLMs due to the size of the dataset and computational constraints.
+Dataset Availability
+We collected data provided by the YouTube API and did not perform additional scraping.
+We only collected public videos and comments, and made no attempt to collect private information.
+We anonymized the author IDs and did not perform any author-level analysis.
+The code used for construction of the dataset and the dataset itself is provided at https://doi.org/10.5281/zenodo.10493803.
+Limitations and Future Work
+Our dataset is limited to the comments from one channel, and the comments we could collect (not all videos posted to the channel allowed comments).
+We hope to expand our dataset to include the comments towards videos posted by armies around the world, to evaluate public opinion towards the army, across the world.
+In addition, our dataset is limited to relatively small LLMs due to computational constraints.
 In the future, we hope to expand the dataset annotation to different types of LLMs with varying prompting schemes to better understand the stance annotation capabilities of LLMs.
-Conclusion
+VI Conclusion
 Understanding the stance expressed in YouTube comments towards an entity is essential for gaining insights into public opinion regarding the authors of the videos.
-In this context, we introduce the DIVERSE dataset, a dataset that captures the stance towards the US Army.
-This dataset is formulated from comments gathered from the U.S.
-Army’s YouTube videos.
+Here, we introduce the DIVERSE dataset, a dataset that captures the stances towards the US Army.
+This dataset is formulated from comments gathered from the US Army’s YouTube videos.
 We constructed the final stance variable utilizing weak labels, including the identification of hate speech and sarcasm, and incorporated stance inferences from Large Language Models.
-By employing a data programming model, we amalgamated these diverse labels to generate the final stance annotations.
+By employing weak supervision, we amalgamated these diverse labels to generate the final stance annotations.
 We anticipate that this dataset will prove invaluable to researchers investigating public sentiments and stances toward military entities.
-Acknowledgments
-The research for this paper was supported in part by the Center for Informed Democracy and Social-cybersecurity (IDeaS) and the Center for Computational Analysis of Social and Organizational Systems (CASOS) at Carnegie Mellon University.
+Acknowledgements
+The research for this paper was supported in part by the Center for Informed Democracy and Social-cybersecurity (IDeaS) at Carnegie Mellon University.
 This work was also conducted within the Cognitive Security Research Lab at the Army Cyber Institute.
@@
 Government.
-Appendix A Appendix 1: Ethics Paper Checklist
-  1. 1.
-    For most authors…
-    1. (a)
-      Would answering this research question advance science without violating social contracts, such as violating privacy norms, perpetuating unfair profiling, exacerbating the socio-economic divide, or implying disrespect to societies or cultures?
-Yes
-    2. (b)
-      Do your main claims in the abstract and introduction accurately reflect the paper’s contributions and scope? Yes
-    3. (c)
-      Do you clarify how the proposed methodological approach is appropriate for the claims made? Yes
-    4. (d)
-      Do you clarify what are possible artifacts in the data used, given population-specific distributions?
-We discuss the limitations of the dataset as well as give basic statistics on the dataset to clarify population-level distributions.
-    5. (e)
-      Did you describe the limitations of your work? Limitations are described in the limitations subsection in the discussion section.
-    6. (f)
-      Did you discuss any potential negative societal impacts of your work? Yes
-    7. (g)
-      Did you discuss any potential misuse of your work? Yes
-    8. (h)
-      Did you describe steps taken to prevent or mitigate potential negative outcomes of the research, such as data and model documentation, data anonymization, responsible release, access control, and the reproducibility of findings?
-Yes.
-We follow standard anonymization procedures for the social media data
-    9. (i)
-      Have you read the ethics review guidelines and ensured that your paper conforms to them? Yes
-  2. 2.
-    Additionally, if your study involves hypotheses testing…
-    1. (a)
-      Did you clearly state the assumptions underlying all theoretical results? NA
-    2. (b)
-      Have you provided justifications for all theoretical results? NA
-    3. (c)
-      Did you discuss competing hypotheses or theories that might challenge or complement your theoretical results? NA
-    4. (d)
-      Have you considered alternative mechanisms or explanations that might account for the same outcomes observed in your study? NA
-    5. (e)
-      Did you address potential biases or limitations in your theoretical framework? NA
-    6. (f)
-      Have you related your theoretical results to the existing literature in social science? NA
-    7. (g)
-      Did you discuss the implications of your theoretical results for policy, practice, or further research in the social science domain? NA
-  3. 3.
-    Additionally, if you are including theoretical proofs…
-    1. (a)
-      Did you state the full set of assumptions of all theoretical results? NA
-    2. (b)
-      Did you include complete proofs of all theoretical results? NA
-  4. 4.
-    Additionally, if you ran machine learning experiments…
-    1. (a)
-      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? We include all of the code in a URL
-    2. (b)
-      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? Yes.
-    3. (c)
-      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? NA
-    4. (d)
-      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? Yes.
-    5. (e)
-      Do you justify how the proposed evaluation is sufficient and appropriate to the claims made? Yes.
-    6. (f)
-      Do you discuss what is “the cost“ of misclassification and fault (in)tolerance? Yes.
-  5. 5.
-    Additionally, if you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
-    1. (a)
-      If your work uses existing assets, did you cite the creators? NA
-    2. (b)
-      Did you mention the license of the assets? NA
-    3. (c)
-      Did you include any new assets in the supplemental material or as a URL? NA
-    4. (d)
-      Did you discuss whether and how consent was obtained from people whose data you’re using/curating?
-NA.
-All data is publicly available data.
-We were also asked to analyze this data by the organization that owns the YouTube channel from which the data comes from.
-    5. (e)
-      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content?
-Yes.
-We mention the data contains various types of conspiracy theories, hate speech ad misinformation.
-We also anonymize the data to remove user names to and within comments
-    6. (f)
-      If you are curating or releasing new datasets, did you discuss how you intend to make your datasets FAIR (see fair)? Yes.
-    7. (g)
-      If you are curating or releasing new datasets, did you create a Datasheet for the Dataset (see gebru2021datasheets)? Yes.
-  6. 6.
-    Additionally, if you used crowdsourcing or conducted research with human subjects…
-    1. (a)
-      Did you include the full text of instructions given to participants and screenshots? NA
-    2. (b)
-      Did you describe any potential participant risks, with mentions of Institutional Review Board (IRB) approvals? NA
-    3. (c)
-      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? NA
-    4. (d)
-      Did you discuss how data is stored, shared, and deidentified? Yes

### v3

@@
 Abstract
-Stance detection of social media text is a key component of many real-world applications like evaluating marketing campaigns, evaluating political policies or candidates, or evaluating information environments.
-However, creating automatic stance labeling systems requires the manual annotation of stances, which is both tedious and resource-intensive.
-This paper introduces an stance labeling method that makes use of weak signals of sentence tone, then consolidating these signals with a Data Programmingmodel for the final stance label.
-In a time of international conflict, understanding the public opinion towards the country’s military is crucial for recruitment.
-We present DIVERSE, a dataset involve stances towards YouTube videos of the US military [^1] Datasetavailableathttps://doi.org/10.5281/zenodo.10493803.
-On average, the videos have 200 comments each, and the stances skew slightly towards the “against” characterization for both the US army and the video.
+Public opinion of military organizations significantly influences their ability to recruit talented individuals.
+As recruitment efforts increasingly extend into digital spaces like social media, it becomes essential to assess the stance of social media users toward online military content.
+However, there is a notable lack of data for analyzing opinions on military recruiting efforts online, compounded by challenges in stance labeling, which is crucial for understanding public perceptions.
+Despite the importance of stance analysis for successful online military recruitment, creating human-annotated, in-domain stance labels is resource-intensive.
+In this paper, we address both the challenges of stance labeling and the scarcity of data on public opinions of online military recruitment by introducing and releasing the DIVERSE dataset:  [^1] https://doi.org/10.5281/zenodo.10493803.
+This dataset comprises all comments from the U.S.
+Army’s official YouTube Channel videos.
+We employed a state-of-the-art weak supervision approach, leveraging large language models to label the stance of each comment toward its respective video and the U.S.
+Army.
+Our findings indicate that the U.S.
+Army’s videos began attracting a significant number of comments post-2021, with the stance distribution generally balanced among supportive, oppositional, and neutral comments, with a slight skew towards oppositional versus supportive comments.
 Index Terms:
@@
 I Introduction
-Today’s world is rife with armed conflict: the Russia-Ukraine war, the Israel-Hamas conflict, the Yemeni Civil War, Somali Civil War… and so forth.
-The US is involved in a number of these conflicts, and also has the mammoth task of defending her own.
-During conflict, it is crucial to know the stance of the people towards its own defending forces, as a gauge of the support the forces have from the public.
-However, US military is facing a recruitment crisis.
-For the last few years, the US Army have struggled to meet recruiting targets to maintain their desired force structure (Brown, 2023).
+The U.S.
+military is facing a recruitment crisis.
+For the last few years, military organizations like the U.S.
+Army have struggled to meet recruiting targets to maintain their desired force structure (Brown, 2023).
 Social media plays an important role in discerning stances (AlDayel and Magdy, 2021), and in recruitment (Ng and Cruickshank, 2023).
-Thus, it is critical to understand how messages from the military is received on social media, for it may be used to shape beliefs of users.
+Thus, it is critical to understand how messages from the military are received on social media, as it can shape recruiting messaging.
 More broadly speaking, the ability to evaluate how social media content is being received by a target audience is crucial for any marketing campaign, be it a war recruiting effort or a organizational recruitment effort or for influencing consumer behavior.
@@
 Several articles have proposed methods and datasets for the task of stance detection (Allaway and McKeown, 2023; Du Bois, 2007).
-However, existing datasets explore stance as expressed on a single social media platform, mainly X, and only have a single stance per comment (Sobhani et al., 2017; Conforti et al., 2020; Mohammad et al., 2016; Hossain et al., 2020; Kochkina et al., 2018; Villa-Cox et al., 2020).
+However, existing datasets explore stance as expressed on a single social media platform, mainly Twitter/X, and only have a single stance per comment (Sobhani et al., 2017; Conforti et al., 2020; Mohammad et al., 2016; Hossain et al., 2020; Kochkina et al., 2018; Villa-Cox et al., 2020).
 Since stance detection remains a difficult task due to the context-dependent nature of the definition of stance (e.g., the requirement of a target for stance to be defined), along with varying definitions of stance (Ng and Carley, 2022), all of the aforementioned stance datasets rely on human annotation to create the stance labels.
 This can result in human errors being present in the datasets as well as greatly reduce the size of the datasets to be in the thousands since human annotation requires a lot of resources and time.
-To address these shortcomings in prior stance dataset, we propose a new stance dataset based on YouTube comments and labeled with machine labeling through the data programming paradigm of weak supervision.
-This dataset is a collection of YouTube comments posted on videos uploaded by the US Army [^2] Offical US Army YouTube Channel: https://www.youtube.com/@usarmy.
-The comments are annotated for weak labels representing tone (hate speech, sarcasm, sentiment), inference of two LLMs (stance towards video, stance towards army), then combining these labels to produce the final stance labels of the comment towards the US Army.
-With this dataset, we develop a automated and systematic approach to labeling stances, aiding in understanding critical issues with the public opinion towards the army and its recruitment efforts.
-Our contributions of this paper are as follows:
-  • The use of an data-programming based stance-labeling methodology.
-We combine recent advances for creating large, high-quality labeled datasets Ratner et al.
-(2016, 2019) with contextual knowledge to generate quality stance labels.
-This involves making use of weak labels such as the presence of hate speech, the sentiment of the content, the presence of sarcasm, known keywords, and the evaluation of Large Language Models (LLMs), which are combined using weak supervision to determine the final stance label.
+To address these shortcomings in prior stance datasets, we propose a new stance dataset based on YouTube comments and labeled with machine-assisted labeling through the data programming paradigm of weak supervision.
+This dataset is a collection of the YouTube comments posted on videos uploaded by the U.S.
+Army [^2] Offical U.S.
+Army YouTube Channel: https://www.youtube.com/@usarmy.
+The comments are first annotated for weak labels representing aspects of stance expression (hate speech, sarcasm, sentiment).
+The comments are then labeled by prompting three different LLMs.
+Finally, these weak labels are combined to produce the final stance labels of the comments towards the U.S.
+Army and its videos.
+With this data set, we can better understand critical issues with the public opinion toward the U.S.
+Army and its recruitment efforts.
+Our contributions in this paper are as follows:
   • A novel benchmark stance dataset of YouTube comments.
@@
 Another difference is that our dataset labels stances towards more than one target in a comment (the only other being Sobhani et al.
-(2017)), and provides labels to weak signals including: the presence of contentious issues, conspiracy theories, and other emotional content.
+(2017)), and provides labels to weak signals including the presence of contentious issues, conspiracy theories, and other emotional content.
 Finally, this dataset has more than three times the number of labeled comments over the current largest benchmark dataset (Conforti et al.
-(2020) at ${}32,000$ data points vs 173,013 data points).
+(2020) at ${}32,000$ data points vs $>$216,000 data points in ours).
+  • The use of a data-programming based stance-labeling methodology.
+We combine recent advances for creating large, high-quality labeled datasets with contextual knowledge to generate quality stance labels Ratner et al.
+(2019); Huang et al.
+(2024); Smith et al.
+(2024).
+This involves making use of weak labels such as the presence of hate speech, the sentiment of the content, the presence of sarcasm, known keywords, and the evaluation of Large Language Models (LLMs), which are combined using weak supervision to determine the final stance label.
 II Related Work
@@
 The incorporation of weak supervision into the realm of stance detection holds the potential to create quality labeled datasets with much less human annotation.
+Finally, it is important to note that some parallel works to ours have explored combining weak supervision with LLMs to create labels.
+In Smith et al.
+(2024), the authors prompt an LLM to provide possible labels for examples and then combine these weak labels through weak supervision to produce final labels, which are significantly better than what an LLM can produce via zero-shot.
+Huang et al.
+(2024) use an LLM to create weak labeling functions for a dataset and then label the dataset with those weak labeling functions and, finally, combine the weak labels through weak supervision to produce a labeled dataset.
+In all cases, the combination of an LLM and weak supervision produces better improved labeling outcomes from a zero-shot scenario.
 III Dataset Creation
-This dataset contains comments of YouTube comments, each labeled for a set of two final stances: (a) the stance of the comment towards its source video, and (b) the stance of the comment towards the US Army.
+This dataset was collected to understand the reception of YouTube users toward the US Army’s marketing material on YouTube, thereby understanding their support/opposition towards the military.
+Each comment is labeled for two stances: (a) the stance of the comment towards its source video, and (b) the stance of the comment towards the US Army.
 There are three possible options for stances: “supports”, “against”, and “neutral”.
-This dataset also contains the weak labels used to fold into the final stance through weak supervision.
+The following parts of this section details the data collection and labeling procedures.
 III-A Data Collection
-This dataset was collected to understand the reception of YouTube users toward the US Army’s marketing material on YouTube, thereby understanding their support/opposition towards the military.
 We collected comments on the videos of the US Army’s official YouTube channel (username: @usarmy [^3] https://www.youtube.com/usarmy) using the YouTube Data API [^4] https://developers.google.com/youtube/v3/docs/comments/list.
 This returned both comments as well as replies to comments.
-Collection took place from 2 to 5 October 2023, where we retrieved a total of 173,013 comments across 858 videos.
+Collection took place from 2 to 5 October 2023 and again from 2-9 April 2024, where we retrieved a total of 216,385 comments across 1,082 videos.
 For each comment, we collected the following information:
@@
   • conversation_id: a unique id of the conversation that a comment belongs to. This is also the id of the root comment
-  • video_id: the unique id of the video the comment is commenting on
-  • name: the title of the video the comment is commenting on
-  • description: a text description of the video the comment is commenting on
+  • video_id: the unique id of the video of the comment
+  • name: the title of the video of the comment
+  • description: a text description of the video of the comment
   • timestamp: the time and date of the publication of the video the comment is commenting on
@@
 These labels were then combined through a weak supervision model to produce the final labels.
+In this way, we include both specialty functions similar to Huang et al.
+(2024) as well as LLM-direct labels in Smith et al.
+(2024) into the weak supervision process.
 Figure 1, depicts the entire data annotation methodology workflow.
 We detail key aspects of the workflow in the following sections.
-Figure 1: Pipeline of stance labeling methodology note: change LLM box to 2 boxes not 3
+Figure 1: Pipeline of stance labeling methodology
 III-C Annotating Weak Labels
-To label the comments for stance, we first produced weak labels.
-To do this we leveraged dataset-specific traits, domain expertise and LLMs.
+To create the weak labels we leveraged dataset-specific traits, domain expertise and LLMs.
 For example, in this dataset, the presence of hate speech or sarcasm frequently indicates the comment being against the video and by extension, the user that uploads it, which is the US Army.
@@
 Several machine learning algorithms ranging from logistic regression to neural network models have been constructed to perform automatic hate speech detection Fortuna and Nunes (2018).
-For this weak label, we used the dehatebert model from HuggingFace [^5] Hate-speech-CNERG/dehatebert-mono-english Aluru et al.
-(2020), a multi-lingual fine-tuned BERT model.
+For this weak label, we used the hate speech classification model from Kralj Novak et al.
+(2022)  [^5] https://huggingface.co/IMSyPP/hate_speech_en.
 This model was used as is without any fine-tuning.
-For each input text, this model returns a binary label of “hate” or “no_hate”, and we considered the “hate” response as an “against” stance label.
+For each input text, this model returns three different types of “hate” label or an “acceptable” label.
+We considered the “hate” response as an “against” stance label, and any other response as an “abstain.”
 III-C2 Sarcasm
@@
 We used a multilingual sarcasm detector666helinivan/multilingual-sarcasm-detector which trains a BERT-based sarcasm detector from newspapers text in English, Dutch and Italian.
-For each input text, this model returns a label of “sarcasm” or “no_sarcasm”, and we considered the “sarcasm” response as an “against” stance label
+For each input text, this model returns a label of “sarcasm” or “no_sarcasm”, and we considered the “sarcasm” response as an “against” stance label, and any other response as an “abstain.”
 III-C3 Sentiment
@@
 For each input text, this model returns a score between 0 and 1 for three possible sentiment labels of “positive”, “negative”, and “neutral”.
-We considered a “positive” sentiment greater than 0.75 as a “supports” stance and a “negative” sentiment greater than 0.75 as a “against” stance and any other value as a “neutral” stance.
+We considered a “positive” sentiment greater than 0.75 as a “supports” stance and a “negative” sentiment greater than 0.75 as a “against” stance and any other value as a “abstain” stance.
 III-C4 Use of Keywords
-We manually determined a set of keywords and phrases that upon present, would represent either one of the stance labels.
+We manually determined a set of keywords and phrases that when present, would represent either one of the non-neutral stance labels.
 For each comment, we identified if any words from the set of keywords were present and assigned the corresponding stance to this weak label.
 For example, keywords and phrases like “hooah”, “awesome”, and “god bless” were associated with a “supports” stance, while “woke”, “murder”, and “propaganda” were associated with an “against” stance.
+the full list of these words is available in our released code [^8] https://doi.org/10.5281/zenodo.10493803.
 III-C5 Stance from Large Language Models
@@
 Due to the size of the dataset, we opted to use open-source LLMs which could be run on local hardware.
-Specifically, we used two different LLMs: (a) Flan-UL2 (Tay et al., 2022), which is an encoder-decoder architecture and (b) Mistral-7B (Jiang et al., 2023), which is a decoder-only architecture.
-We used the Flan-UL2 model in a zero-shot setting (i.e., we did not perform any fine-tuning on the model) (Ziems et al., 2023).
-For the Mistral-7B model, we fine-tuned the model on a collection of benchmark stance datasets, including Semeval2016 (Mohammad et al., 2016), Election2016 (Sobhani et al., 2017), srq (Villa-Cox et al., 2020), wtwt (Conforti et al., 2020), phemerumors (Kochkina et al., 2018), and covid-lies (Hossain et al., 2020).
+Specifically, we used three different LLMs: (a) Flan-UL2 (Tay et al., 2022), which is an encoder-decoder architecture, (b) Mistral-7B (Jiang et al., 2023), which is a decoder-only architecture, and (c) Llama-3-8B Dubey et al.
+(2024), which is another decoder-only architecture.
+We used the Flan-UL2 and Llama-3 models in a zero-shot setting only (i.e., we did not perform any fine-tuning on the model) (Ziems et al., 2023).
+For the Mistral-7B model, we both used it in a zero-shot setting and we fine-tuned the model on a collection of benchmark stance datasets, including Semeval2016 (Mohammad et al., 2016), Election2016 (Sobhani et al., 2017), srq (Villa-Cox et al., 2020), wtwt (Conforti et al., 2020), phemerumors (Kochkina et al., 2018), and covid-lies (Hossain et al., 2020).
 Based on the research in (Ng and Carley, 2022), which found that the inclusion of more benchmark datasets into training improved the generalizability of the trained model, hence we trained on all of the datasets.
@@
 The model was trained for two epochs using an AdamW optimizer (Loshchilov and Hutter, 2017), a learning rate of $41e-$ 4, and a linear learning rate scheduler with 50 warm-up steps.
-To feed the data to the model during training, we adopted a prompt, with two variations in the task portion of the prompt, that included context about stance classification, as has been done in zero-shot prompting for stance in other works.
+To feed the data to the model during training, we adopted a prompt that included context about stance classification, as has been done in zero-shot prompting for stance in other works.
 The following was the training prompt:
@@
 Where the [dataset entity] was varied depending on the benchmark dataset (e.g., for Election2016, it was ‘politician’, while for phemerumors it was ‘rumor’).
-For the other variation of the prompt, the “Analyze…” sentence was replaced with a question-based tasking in the form of: “What is the stance of the following social media statement towards the provided [dataset entity]”.
 For the srq dataset, we changed the prompt slightly to incorporate the post that the reply is replying to as the task in that dataset is to label the stance of the reply to the post, versus the stance of the post about a topic or person.
@@
 The training of the model was done on a local computer running Ubuntu Linux with a 96-core processor and three A6000 GPUs.
-Overall, this resulted in three, different LLMs for labeling the comments: one off-the-shelf UL2 model and two, fine-tuned Mistral models.
+Overall, this resulted in four, different LLMs for labeling the comments: one off-the-shelf UL2 model, one off-the-shelf Llama-3 model and two Mistral models.
+Finally, we also fine-tuned a small BERT model as was done in Ng and Carley (2022) on all of the benchmark datasets previously described and labeled the stances with that fine-tuned model as well.
 Analyzing stances with LLMs
-In analyzing the comments for stances using LLMs, we used two different prompts, one that only included the comment, and one that included the comment that a comment was replying to, if it was replying to a comment, with two different variations on each prompt.
+In analyzing the comments for stances using LLMs, we used two different prompts, one that only included the comment, and one that included the comment that a comment was replying to, if it was replying to a comment.
 The labeling prompt without incorporating the reply structure is:
@@
 stance:
-As with the training prompt, the variations for each prompt came from changing the task part of the prompt (i.e., the “Analyze…”) sentence to a question-based task prompt.
-These variations were only used with the Mistral model fine-tuned by the question prompt Thus, each comment was labeled for a stance by an LLM six different times: once with the reply prompt and once without the reply prompt, for the UL2 model, the stance-tuned Mistral model, and the stance-tuned Mistral model with the question-task prompt.
-Since we are labeling the comment for its stance toward the US Army as well as the video it is commenting on, this results in a total of 12 LLM labeling runs across the entire dataset.
+We also experimented with a question answering phrasing of the task during labeling, but found the analyze phrasing to generally perform slightly better in benchmark datasets.
+Overall, each comment was labeled for a stance by an LLM seven different times: once with the reply prompt and once without the reply prompt for the UL2 model, the Llama-3 model, and other Mistral model, and with the without the reply prompt by the stance-tuned Mistral model.
+Since we are labeling the comment for its stance toward the US Army as well as the video it is commenting on, this results in a total of 14 LLM labeling runs across the entire dataset.
+To combine the LLM-derived labels we choose to ensemble the LLM labels together using majority vote.
+Specifically, we considered those labels where more than half (i.e., 4 or more of the LLM labelers) agreed on a stance for a comment as producing a stance, otherwise we assigned the label from the LLMs as “abstain.” Although this minorly reduced some of the coverage of the LLM labels, it also improved the quality of the LLM labels overall.
 III-C6 Use of Analogous Stances
-Finally, in the case of labeling the stance of the comment toward its respective video, we also used an analogous stance for a weak label.
-We used the stance of the comment toward the US Army.
-Generally speaking, if a comment was against the US Army, it would also be against a video posted by the US Army, and the use of complementary stances has been shown to improve performance in the stance detection task Sobhani et al.
+Finally, in the case of labeling the stance of the comment toward its respective video, we also used an analogous stance for a weak label; we used the stance of the comment toward the U.S.
+Army as a weak label for the stance toward the video.
+Generally speaking, if a comment was against the U.S.
+Army, it would also be against a video posted by the U.S.
+Army, and the use of complementary stances has been shown to improve performance in the stance detection task Sobhani et al.
 (2017); Zhang et al.
 (2023b).
-Thus, we used the final stance toward the US Army produced by weak supervision on the aforementioned weak labels as a weak label for the stance toward the video.
-It is important to note that this relationship does not work the other way; we found several examples where commentators would express a supportive stance toward the US Army but would be very against a particular video.
+Thus, we used the final stance toward the U.S.
+Army produced by weak supervision on the aforementioned weak labels as a weak label for the stance toward the video.
+It is important to note that this relationship does not work the other way; we found several examples where commentators would express a supportive stance toward the U.S.
+Army but would be very against a particular video.
 III-D Final Stance Determination
@@
 III-E Human evaluation
-After machine annotation, we had four individuals annotate how many?
-what percentage of total?
-comments.
-For each comment, the authors independently annotated the stance towards the US Army and the stance towards the video.
-The manual annotations were then analyzed for inter-annotator agreement, and the machine annotations were compared against the manual ones.
+In order to assess the quality of the machine-produced labels, we needed to validate these labels by humans.
+Since it is impractical to label the entire dataset by humans, we randomly selected 1,000 comments which were labeled by four annotators.
+We then assessed the label agreement among the annotators and combined the human-labels by majority vote.
+Using these labels, we assessed the quality of the machine-produced labels.
 IV Data Properties and Analysis
 This section describes some exploratory data analysis performed on this dataset.
-The dataset consists of 858 videos published from 1 June 2010 through 5 October 2023 (final date of collection) on the US Army’s official YouTube Channel.
+The dataset consists of 1082 videos published from 1 June 2010 through 1 April 2024 on the US Army’s official YouTube Channel.
+From these videos, there are a total of 215,509 comments.
 IV-A Dataset Properties
@@
 Word-level Statistics of Comments Value
-Mean 16.77
-Standard Deviation 32.67
+Mean 16.66
+Standard Deviation 33.13
 Minimum 1.00
@@
 Statistics of comments per video Value
-Mean 201.65
-Standard Deviation 873.19
+Mean 199.17
+Standard Deviation 883.41
 Minimum 1
-25th Percentile 5
-50th Percentile (Median) 27
-75th Percentile 77
-Maximum 16289
+25th Percentile 7
+50th Percentile (Median) 37
+75th Percentile 86
+Maximum 18999
 Mode 1
@@
 Table III summarizes the label counts for each of the stances.
-Interestingly, from the label counts we also observe both a relatively even distribution of labels across the stances but also the “neutral” stance being the minority label.
-In most other benchmark stance datasets, there is a strong label skew, particularly for datasets like Hossain et al.
+From the labeling we can see some label skew.
+First, we observe a label skew toward the “neutral” label class, which is characteristic of all other stance-labeled datasets (e.g., Hossain et al.
 (2020) and Kochkina et al.
-(2018), the “neutral” stance tends to be the majority stance label.
-Since this is a first-of-its-kind dataset for both the targets as well as the data domain (i.e., YouTube) it’s not clear if the targets tend to inspire stronger stance-taking behavior in comments or if the domain of YouTube causes stronger stance-taking behavior.
-Regardless, upon manual inspection of some of the labels, we do find that these stance labels are accurate and that there is a lot of strong stance-taking within the comments.
+(2018)).
+We do also, however, observe a slight label skew toward “against” stances versus “supportive” stances.
+Since these are comments meant to advocate for a brand (i.e., the U.S.
+Army), the skew in valenced stances against the U.S.
+Army and its videos is a cause for concern.
+Finally, we also observe that through the use of the weak labels, we were unable to label to a couple of thousand data points (around $1\%$ of the data) for each stance.
 Stance U.S. Army Video
-against 63676 (34.87%) 67405 (36.92%)
-neutral 49153 (26.91%) 49310 (27.01%)
-supports 60183 (32.99%) 56297 (30.88%)
+abstain 2373 (1.10%) 1564 (0.73%)
+against 49817 (23.12%) 63222 (29.34%)
+neutral 122028 (56.61%) 102479 (47.56%)
+supports 41291 (19.16%) 48244 (22.37%)
 TABLE III: Labels Counts (and percentages of the total number) for Each of the Stance Targets
@@
 We further observe that the two prominent periods of comments do differ somewhat in their compositions of stance labels.
-In the first peak of comments, centering in September of 2022, we see more “against” stance labels than “supports” or “neutral”.
-Whereas in the second spike of comments around May-June of 2023, we see more “supports” comments than “against” or “neutral”.
+In the first peak of comments, centering in September of 2022, we see more “against” stance labels than “supports”.
+Whereas in the second spike of comments around May-June of 2023, we see more “supports” comments than “against”.
 Thus, overall, there is a moderate balance among the stances expressed in the comments toward the videos, but this balance does shift over time and between videos.
@@
 To better understand these differences in stances, we looked at the videos with the most positive, neutral and negative receptions.
-Among the videos with more than 100 comments, we found the video, “Honoring Our Fallen Heroes - Taps” had the highest percentage of supportive comments at 76.5% of the comments supporting the video.
-On the opposite end was the video “What does the Army of the future look like?” which features a short interview with former Sergeant Major Grinston on the future of the US Army.
-The video had 66.9% of comments that were against it.
-And, the most neutrally received video was “Army #PopQuiz - What training is this?” which received 66.5% of its comments as being neutral toward the video, many of them answering the question posed in the video.
+Among the videos with more than 100 comments, we found the video, “Honoring Our Fallen Heroes - Taps” had the highest percentage of supportive comments at 61.5% of the comments supporting the video.
+On the opposite end was the video “Sergeant Major of the Army Grinston addresses the report of the Fort Hood Independent Review” which features a short interview with former Sergeant Major Grinston on the investigation at fort Hood following the murder of the Venessa Guillen.
+The video had 55.9% of comments that were against it.
+And, the most neutrally received video was “Army #PopQuiz - What training is this?” which received 81.6% of its comments as being neutral toward the video, many of them answering the question posed in the video.
 Finally, since this dataset was labeled with more than one stance, we can also look at the relationships between the two stances.
@@
 IV-C Comparison with human evaluation
-The inter-annotator agreement was Iain, results here. Comparing the machine-based annotations with the human-based annotations, we get an accuracy of how many. elaborate on accuracy
-IV-D Leave-One-Out Stance Detection
-To further validate the stance labels produced by the proposed weak supervision method, we performed a leave-one-out, out-of-domain experiment referenced from (Ng and Carley, 2022).
-We trained a model on the stance labels and then evaluated that model on the other benchmark datasets.
-In our test, we trained a cased BERT model [^9] https://huggingface.co/bert-base-cased (Devlin et al., 2018) on only the stance labels from the DISMISS dataset and then evaluated the model on all of the benchmark datasets.
-The model was trained for 3 epochs using a learning rate of $41e-$ 4 with linear decay, a batch size of 32, and was trained to classify text into three different labels.
-The text that was input to the model was given as:
-“target: {target of stance} [SEP] comment: {comment}”
-In this test, if the proposed data labels are capturing the stances correctly, then we should see similar performance on the benchmark datasets when predicting stance, as was seen in Ng and Carley (2022).
-Table IV has the out-of-domain test results.
-The results are evaluated using the Macro F1 score as a suitable evaluation method due to the almost equal distribution of stance labels.
-On average, the model trained on stances only from this dataset achieves a macro F1 score of 0.3043.
-This is similar to the average accuracy of 0.303 in a past work that performed leave-one-out stance detection using multiple BERT-based classifiers Ng and Carley (2022).
-This therefore suggests that the stance labels proposed in this paper are valid stance labels.
-Dataset Macro F1 Score
-semeval2016 0.3517
-election2016 0.3439
-covid-lies 0.3020
-phemerumors 0.2295
-wtwt 0.3018
-srq 0.2971
-Average 0.3043
-TABLE IV: Macro F1 Scores for Model Applied to Different Datasets
+Finally, we analyzed how well machine labeling, using state–of–the–art techniques, of stance, performed.
+To do this, we created a validation dataset as previously described and had four human labelers label that data set.
+For the stance toward the U.S.
+Army, we had an average Cohen’s Kappa McHugh (2012) across all of the human labelers of $0.863$ and for the stance toward the videos of $0.797$.
+Thus, we were able to obtain high agreement between the labelers on stance of the comments (note that we also released this labeled validation dataset).
+Compared to the output from the weak supervision process, we obtained moderately successful machine labels.
+For the stance toward the U.S.
+Army of the comments, the accuracy and Micro F1-scores were $0.68$ and for the stance toward the videos of the comments, the accuracy and Micro F1-scores were $0.639$.
+When we similarly compared the results of each of the LLMs in the labeling task (see Table IV), we do see an improvement of the weak supervision-produced labels over the best-performing LLMs labels, especially for the stance toward the videos.
+Labeling Model U.S. Army Video
+mistral zero-shot with comments 66.90% 56.50%
+llama-3 zero-shot with comments 41.40% 50.90%
+mistral zero-shot 64.40% 55.10%
+llama-3 zero-shot 41.20% 40.20%
+ul2 zero-shot with comments 53.00% 45.00%
+ul2 zero-shot 49.00% 45.80%
+mistral adapter 49.70% 26.50%
+small model 52.50% 32.30%
+TABLE IV: Micro-F1 scores for U.S. Army and Video stances, by different models and prompts, on the validation dataset.
+From the analysis of the LLM labelers, we notice that there are distinct differences in performance between models and prompting schemes, with the Mistral LLM Jiang et al.
+(2023) and the prompt scheme incorporating replies in the comments generally performing better.
+To further understand these results, we plotted the confusion matrix between the human-derived and machine-derived labels in Figure 6.
+From these matrices, we can observe that for valenced stances, the machine-label stances generally match those from human-annotation; much of loss of accuracy in the machine-labeled stances comes from mis-labeling the neutrally-stanced comments.
+Upon inspection of some these examples, it seems there are a number of highly-positive sounding comments that are unrelated to the video or the Army as well as some comments where sarcasm or coded language (e.g., ‘goy’, ‘zog’ were missed by the LLMs.
+Figure 6: Confusion matrices between the machine-derived labels and the human-derived labels for the validation set.
 V Discussion
-Public opinion towards a country’s army is crucial, as it signals the level of trust in the army and the willingness of citizens to join the forces.
-A positive opinion facilitates army recruitment and support, especially in times of crisis.
+Public opinion towards a country’s military is crucial, as it signals the level of trust in the government writ large and the willingness of citizens to join the military.
+A positive opinion facilitates military recruitment and support, especially in times of crisis.
 Social media engagement plays a pivotal role in shaping this opinion.
-This paper introduces a dataset that supports such understanding and associated analyses, comprising comments on the US Army’s official YouTube videos annotated for their stance towards the Army and the video content itself.
+This paper introduces a dataset that supports such understanding and associated analyses, comprising comments on the US Army’s official YouTube videos annotated for their stance towards the Army and the video itself.
 To construct this dataset, we employed weak supervision, which leverages weak labels for automated labeling.
@@
 The combination of dataset-specific weak signals and the linguistic understanding power of LLMs produces high-quality labels for stance in a more scalable manner compared to purely manual or machine-generated labels.
-Finally, our Leave-One-Out experiment produces a standard macro F1 score of around 0.34, a score that is similar to past work.
-This suggests that the stance labels produced by the proposed methodology are indeed valid.
-This result also highlights findings from past work that trained stance detection models do not necessarily generalize to other data domains (e.g., YouTube versus Twitter, or targets of videos versus political figures, etc.).
-Thus, we suggest training a stance classification model using multiple benchmark datasets so the model can better recognize the nuances of the stances and targets involved.
-The low F1 score also indicates the difficulty of the stance classification task, whereby language models do struggle to differentiate the nuances in opinions towards a target entity, despite having helper annotations like hate and sentiment to provide indications of stances.
+In this work we combined the existing state-of-the-art methods for automated labeling of a dataset on the difficult task of stance labeling.
+We found that these approaches were able to produce reasonable labels, relative to a human-labeled validation dataset, but also that there is still room for improvement in the development of automated labeling techniques, especially for challenging labeling tasks, like stance.
 This, therefore, opens up avenues for exploration of improving the stance classification methods, especially on out-of-domain tasks, to provide a more accurate label towards an entity and therefore more accurate downstream insights.
@@
 Beyond its primary application in training and evaluating new stance detection techniques and models, the dataset introduces distinctive traits not present in current benchmarks.
-Notably, it originates from YouTube, offering a departure from the prevalent X-based datasets.
+Notably, it originates from YouTube, offering a departure from the prevalent X or Twitter-based datasets.
 Furthermore, the dataset features multiple stance targets and exhibits a substantial degree of stance-taking behavior and controversial content (i.e., a high proportion of “supports” and “against” stances).